### v2

# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co-evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value-function loss. Building on these approaches, we introduce the transition prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called co-learnability. By combining these two measures, we present Transition-aware Regret Approximation with Co-learnability for Environment Design (TRACED). Empirical evaluations show that TRACED yields curricula that improve zero-shot generalization across multiple benchmarks while requiring up to 2$\times$ fewer environment interactions than strong baselines. Ablation studies confirm that the transition prediction error drives rapid complexity ramp-up and that co-learnability delivers additional gains when paired with the transition prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED.
1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics [1]. Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand-crafting a training distribution that captures all real-world variability is intractable, and agents often overfit even large training sets, performing poorly out-of-distribution [2, 3].
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent [4]. A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design [4, 5, 6, 7, 8, 9, 10]. Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) [11]. In this paper, we refine regret estimation by explicitly decomposing it into (i) a value error, captured by PVL, and (ii) a transition prediction error, estimated via a transition dynamics model.
We further introduce a metric called Co-Learnability to quantify how training on one task benefits others. For instance, consider three 100-word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy [12, 13, 14], reflecting high Co-Learnability; in contrast, Japanese is typologically distant [15], so gains from Japanese may transfer unefficiently to English, reflecting low Co-Learnability. We present a lightweight estimator of Co-Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
Figure 1: Task Priority Landscape. Task in the top-right quadrant (high difficulty and high co-learnability) are scheduled with the highest priority in the curriculum.
As illustrated in Figure 1, these two ingredients, refined regret and Co-Learnability, are fused into a unified Task Priority score via a rank-based transform (cf. [7]), which governs both new task generation and prioritized replay. We call our algorithm TRACED (Transition-aware Regret Approximation with Co-learnability for Environment Design). We evaluate TRACED on procedurally generated MiniGrid (MG) [16] and BipedalWalker (BW) [17]. In MG, we compare against Domain Randomization (DR) [18], PLR [5], ADD [19], and ACCEL [7], the strongest of these baselines. In BW, we additionally include the state-of-the-art (sota) CENIE [20].
On 12 held-out MG mazes, TRACED achieves a mean solved rate of 66% after 10k updates, outperforming ACCEL’s 60% after 20k updates. In the extreme maze variants PerfectMaze, TRACED reaches 27% success on PerfectMazeLarge and 10% on PerfectMazeXL, representing 35-40% relative gains over ACCEL with only half the training time. In BW, TRACED achieves an 18% improvement over CENIE while halving the wall-clock time. Ablation studies reveal that the transition-prediction error term drives rapid complexity ramp-up, and that Co-Learnability provides additional gains when combined with our regret term. Collectively, these results pave a new path toward sample-efficient curriculum design in UED by combining refined regret approximation with explicit modeling of task relationships.
We provide the full implementation of TRACED to facilitate the reproduction of our main results at https://github.com/Cho-Geonwoo/TRACED.
2 Preliminaries
2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al. [4]. A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory-conditioned policy $\pi\bigl{(}a_{t}|o_{0},a_{0},\dots,o_{t}\bigr{)}$. For a fixed level $\theta$, the utility of policy $\pi$ is the expected discounted return $U_{\theta}(\pi)=\mathbb{E}\!\Bigl{[}\sum_{t=0}^{T}\gamma^{t}\,r_{t}\Bigr{]},r_{t}=\mathcal{R}(s_{t},a_{t})$ with the expectation taken over both the stochastic dynamics and the policy’s choices. We denote an optimal policy on level $\theta$ by $\pi^{\star}_{\theta}\in\arg\max_{\pi}U_{\theta}(\pi)$.
2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ [4]. Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning [5, 7, 20]. The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR [6] evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)-based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
$\mathrm{PVL}(\tau)=\frac{1}{T}\sum_{t=0}^{T}\max({\sum_{k=t}^{T}(\gamma\lambda)^{\,k-t}\,\delta_{k}},0\Bigr{)}$ (1)
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one-step TD-error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one-step TD error, GAE, policy min-margin, and policy least-confidence have also been evaluated in curriculum learning contexts [5].
3 Method
At each step, curriculum design aims to select the tasks that will most effectively drive the agent’s learning progress. Our approach enhances UED by combining (1) a refined regret approximation integrating both value and transition errors and (2) a simple yet principled measure of task co-learnability. We compute a unified Task Priority score from these two quantities, which then governs both new task generation and replay sampling in the UED loop.
3.1 Regret Approximation via Transition Prediction Loss
Since regret quantifies the difficulty an agent experiences on a task, improving regret approximation can lead to more accurate task difficulty estimation and thus more effective curricula.
To address this, we decompose one-step regret at state–action pair $(s,a)$ into two interpretable terms:
$\mathrm{Regret}(s,a)$ $=Q^{*}(s,a)-Q(s,a)$ (2)
$=\gamma\Bigl{(}\mathbb{E}_{s_{1}\sim P(\cdot\mid s,a)}\bigl{[}Q^{*}(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}(\cdot\mid s,a)}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}\Bigr{)}$
$=\gamma\Bigl{(}\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}\,Q^{*}(s_{1},a^{\prime})-Q(s_{1},a^{\prime})\,\bigr{]}}_{\text{(i) Value Error}}+\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}Q(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}}_{\text{(ii) Transition Prediction Error}}\Bigr{)}$
Term (i), the value error, measures the expected discrepancy between the true optimal action-value $Q^{*}$ and the agent’s learned $Q$. Term (ii), the transition prediction error, quantifies the bias introduced by replacing the true transition kernel $P$ with an empirical approximation $\hat{P}$, where $\hat{P}(s^{\prime}\!\mid s,a)$ is the conditional distribution estimated from the agent’s collected experience $\{(s,a,s^{\prime})\}$. Accurately estimating both terms is crucial: term (i) reflects the agent’s value-approximation quality, whereas term (ii) captures how model-mismatch in the dynamics affects long-term returns.
Because $Q^{*}$ and $P$ are unknown, we approximate (i) via the Positive Value Loss (PVL) and (ii) via a learned transition dynamics estimator $f_{\phi}$. Concretely:
Definition 1 (Average Transition Prediction Loss). Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one-step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ (details in Appendix F.3). Over an episode $\tau=(s_{0},a_{0},\dots,s_{T})$, define
$\mathrm{ATPL}(\tau)=\frac{1}{T}\sum_{t=0}^{T}L_{\mathrm{trans}}(s_{t},a_{t}).$ (3)
Combine the two estimates into a single scalar:
$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau),$ (4)
where $\alpha>0$ balances value versus transition terms (see Appendix B).
By explicitly incorporating both components, our regret approximation more faithfully reflects task difficulty and yields more effective curricula. After computing the approximated regret $\widehat{\mathrm{Regret}}(\tau)$ for each sampled task instance, we append this value to the Task Difficulty Buffer (TDB). Consequently, the number of entries in the TDB corresponding to task $i$ exactly equals the number of times task $i$ has been sampled (see Figure 2 for an overview).
Figure 2: Task Difficulty Calculation Workflow. The Task Difficulty Buffer (TDB) records each task’s history of approximated regret. The agent interacts with sampled tasks to collect episode trajectories, which are stored in the Rollouts Buffer (RB). For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition-Prediction Loss (ATPL). Their sum produces the updated task difficulty (approximated regret), which is appended to the TDB to refresh each sampled task’s stored difficulty.
3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step. Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co-Learnability, and combine them into a single Task Priority score.
Definition 2 (Task Difficulty). Let $s_{i}(t)\;=\;\max\{\,s\leq t:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $t$ when task $i$ was sampled (and its approximated regret stored). Then we define
$\mathrm{TaskDifficulty}(i,t)=\begin{cases}\mathrm{TDB}\bigl{(}i,\,s_{i}(t)\bigr{)},&\text{if }s_{i}(t)\text{ is finite},\\[4.0pt] 0,&\text{if }i\text{ has never been sampled before }t.\end{cases}$ (5)
This ensures that $\mathrm{TaskDifficulty}(i,t)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
Definition 3 (Co-Learnability). Beyond difficulty, we wish to capture how training on one task accelerates progress on others. Let $\mathcal{T}_{t+1}$ be the set of tasks replayed at time $t+1$. We define
$\mathrm{CoLearnability}(i,t)=\frac{1}{|\mathcal{T}_{t+1}|}\sum_{j\in\mathcal{T}_{t+1}}\bigl{[}\mathrm{TaskDifficulty}(j,t)-\mathrm{TaskDifficulty}(j,t+1)\bigr{]}$ (6)
i.e.,  the average reduction in difficulty of the replayed tasks induced by sampling $i$. Positive co-learnability indicates that visiting $i$ yields broad transfer benefits.
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proposition 1 confirms that our co-learnability definition precisely measures the mean one-step change in task difficulties, conditional on having replayed task i, when the next sampled task set covers the entire buffer.
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score, then pass it through a rank function, which replaces raw values by their relative ordering, to obtain robust sampling weights as in ACCEL [7]:
$\mathrm{TaskPriority}(i,t)=\mathrm{Rank}\Bigl{(}\mathrm{TaskDifficulty}(i,t)+\beta\,\mathrm{CoLearnability}(i,t)\Bigr{)}$ (7)
where $\beta>0$ trades off challenge versus transfer. We sample tasks for replay proportionally to their computed priority scores, so that higher-priority tasks are more likely to be selected at each step. By ranking, we prevent outliers from dominating the curriculum and ensure a smooth sampling distribution.
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
This guarantees that no task remains “stale” indefinitely: its expected waiting time scales inversely with its initial priority. While the boundedness and non-increasing requirements in Theorem 3.2 may initially seem restrictive, they follow naturally from the agent’s continual learning process. Indeed, Figure 10 shows an overall upward trend in returns during training, indicating that the agent’s performance steadily improves and, consequently, task priority scores tend to decrease over time.
3.3 Overall UED Workflow
Figure 3: Method Workflow Overview. The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer in proportion to their priority scores. (2) Priority Update: we recompute each level’s priority based on our task priority definition (Eq. 7). (3) Task Mutation: the lowest-priority levels are mutated into new variants and reinserted into the buffer.
The overall UED algorithm follows the same loop as ACCEL [7], with the sole modification that task scoring uses our task priority definition (Eq. 7) in place of positive value loss alone. We maintain a policy $\pi_{\phi}$, a level generator G, a Task Buffer $\Lambda$, and the sampling distribution $P_{D}$ over the decision variable $d\in\{0,1\}$, where $d=0$ indicates generator sampling and $d=1$ indicates buffer replay.
At each iteration $t$, we draw $d_{t}\sim P_{D}$ and proceed as follows. If $d_{t}=0$, we sample a fresh task parameter $\theta$ from the generator and run the current policy to collect a trajectory $\tau$ on $\theta$. We then compute the approximated regret score $\widehat{\mathrm{Regret}}(\tau)$ using Eq. 4. If $\widehat{\mathrm{Regret}}(\tau)$ exceeds a predefined threshold, $\theta$ is appended to $\Lambda$.
If instead $d_{t}=1$, we sample a batch of $B$ tasks $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ according to their TaskPriority distribution (Eq. 7). For each $\theta_{k}$, we execute the policy $\pi_{\phi}$ to collect trajectory $\tau_{k}$, perform the standard RL update on $\phi$, and compute its approximated regret $\widehat{\mathrm{Regret}}(\tau_{k})$. We then select the $N_{\text{mutate}}$ tasks with the lowest regret scores and feed each through the task editor to obtain a mutated variant $\tilde{\theta}$. Finally, we replace the corresponding entries in $\Lambda$ with these $\tilde{\theta}$, ensuring that easier tasks are systematically evolved into more challenging ones.
Following each update to the Task Difficulty Buffer, we recompute the $\mathrm{TaskDifficulty}(i,t)$ for every modified task $i$ and update the $\mathrm{CoLearnability}(j,t-1)$ for all previously sampled tasks $j$. These refreshed scores determine the priorities for the next replay phase. The algorithm thus alternates between sampling new tasks and performing prioritized replay until the policy performance converges. The full procedure is detailed in Algorithm 1.
4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE [20]. Detailed descriptions of all baselines can be found in Appendix M.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero-shot test performance on held-out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library [21]. The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance.
All methods use Proximal Policy Optimization (PPO) [22] as the student agent. Following prior work [7, 20], we plot the performance against the number of PPO updates; the corresponding environment interactions per update are reported in Appendix C. For more details on hyperparameters and network structures, refer to Appendix F.
SixteenRooms
Maze
Labyrinth
LargeCorridor
(a) Maze
Stairs
Roughness
(b) BipedalWalker
Figure 5: Held-out Evaluation Environments. (a) Example held-out mazes for zero-shot evaluation: 4 tasks are shown (see Appendix F.1 for all 12 task definitions). (b) Example held-out BipedalWalker terrains for zero-shot evaluation: 2 tasks are shown (see Appendix F.2 for all 6 task definitions).
4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid [16], as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147-dimensional pixel observation and is trained for up to 20k PPO updates.
Figure 6: Zero-Shot Transfer Performance in MiniGrid. Aggregated solved rates on held-out MiniGrid mazes after 10k and 20k PPO updates. TRACED at 10k updates (red) matches or outperforms DR (orange), $\text{PLR}^{\perp}$ (green), ADD (gray), and ACCEL (blue) at 20k updates. Averaged over 10 random seeds.
As shown in Figure 6, our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines. In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and co-learnability scoring. The mean performance mirrors this ordering, showing that even lower-performing seeds see substantial gains. Furthermore, TRACED at 10k produces the smallest optimality gap, closing the distance to an ideal target success rate faster than any competitor; in contrast, ACCEL requires only 20k updates to halve its gap.
Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain for each algorithm at 10k and 20k PPO updates.
TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
13.78 $\pm$ 0.36 12.94 $\pm$ 0.66 26.58 $\pm$ 0.76 22.48 $\pm$ 0.27 45.08 $\pm$ 0.29 14.87 $\pm$ 0.62 31.83 $\pm$ 1.36 5.82 $\pm$ 0.12 12.41 $\pm$ 0.18
Compared to ACCEL, TRACED halves the wall-clock training time while maintaining equivalent or better transfer performance (Table 1). Even relative to ACCEL at the same 10k update budget, TRACED incurs a 6% computational overhead yet delivers a 22% relative increase in mean solved rate. Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero-shot transfer across diverse MiniGrid mazes. Detailed per-task zero-shot results are provided in Appendix E, and the number of environment interactions per PPO update for each method is in Appendix C.
(a) Shortest Path Length
(b) Number of Blocks
Figure 7: Emergent maze complexity metrics. Shortest path length and number of blocks both grow faster under TRACED (pink) than ACCEL (blue). This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning. Averaged over 10 random seeds with shaded 95% confidence intervals.
To analyze emergent environment complexity, we track two structural metrics for each generated maze: the length of the shortest solution path and the number of obstacles (blocks), as plotted in Figure 7. Both metrics rise much more rapidly under TRACED (red) than under ACCEL (blue), showing that our priority scoring more effectively distinguishes easy versus challenging tasks to produce a steadily escalating curriculum. Notably, the ACCEL + ATPL (brown) also drives complexity upward far faster than ACCEL. It closely parallels TRACED’s gains, demonstrating that the transition-prediction component alone contributes greatly to complexity ramp-up.
(a) PerfectMazeLarge
(b) PerfectMazeXL
Figure 8: PerfectMaze Evaluation. (a), (b) Two held-out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero-shot testing. (c) Zero-shot solved rates after 10k and 20k PPO updates (mean $\pm$ std over 10 random seeds). TRACED achieves the highest 10k performance on PerfectMazeLarge and closely matches the best 20k performance on PerfectMazeXL.
To stress-test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks. PerfectMazeLarge consists of $51\times 51$ grids with a maximum episode length exceeding 5k steps, while PerfectMazeXL scales this to $100\times 100$ grids. Figure 8 shows representative levels from each variant. We evaluate zero-shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total). As depicted in Figure 8, on PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$. On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes. Detailed per-baseline results for both benchmarks are provided in Appendix E.
4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous-control BipedalWalkerHardcore environment from OpenAI Gym [23], as modified by Wang et al. (2019) [24]. This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness.
Figure 9: Aggregate zero-shot performance on BipedalWalker terrains. All scores are normalized by the maximum return of 300 in the BipedalWalker domain. TRACED at 10k updates (pink) matches or exceeds ACCEL-CENIE at 20k updates (purple) across all metrics. Averaged over 5 random seeds.
(a) Basic
(b) Hardcore
(c) Stair
(d) Pitgap
(e) Stump
(f) Roughness
Figure 10: Return progression on BipedalWalker terrains. TRACED (pink) consistently outperforms DR (orange), PLR (green), ADD (gray), and ACCEL (blue). Averaged over 10 random seeds with shaded 95% confidence intervals.
Our student agent is trained for 10k PPO updates, while baselines run up to 20k updates. We evaluate zero-shot transfer on six held-out test terrains over 100 episodes each. Figure 9 reveals consistent gains across all four aggregate metrics, median, IQM, mean, and optimality gap, on the six held-out BipedalWalker terrains. Specifically, after only 10k updates, TRACED outperforms all baselines at 20k updates. Although its variance is slightly higher, TRACED’s 10k-update curve matches or exceeds ACCEL-CENIE at 20k, demonstrating both superior sample efficiency and robust generalization. Moreover, as in MG, TRACED reaches the highest performance in half the wall-clock time on BW relative to ACCEL (see Appendix C). These results confirm that our curriculum design remains effective not only in discrete grid worlds but also in complex continuous domains. Refer to Appendix E for detailed results broken down by each terrain.
As shown in Figure 10, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains. In the first 5k updates, TRACED trails ADD, DR, and PLR because it builds upon ACCEL’s seed curriculum, mutating the easiest environments first, before increasing difficulty. After 5k updates, however, TRACED continues to climb steadily, while ADD, DR, and PLR plateau in a stale regime. This sustained upward trajectory demonstrates that TRACED effectively escalates task complexity in tandem with the agent’s improving capabilities, yielding superior long-term performance.
4.3 Ablation Study
Figure 11: Both ATPL and CL are important design choices. Showing per-task zero-shot solved rates for 4 variants: full TRACED (pink), ATPL only (brown), Co-Learnability only (teal), and ACCEL (blue). Averaged over 5 random seeds.
To isolate the contributions, we compare TRACED at 10k updates against three ablations (Figure 11): ATPL only (TRACED – CL), Co-Learnability only (TRACED – ATPL), and the ACCEL baseline (TRACED – ATPL – CL). TRACED (pink) encloses the largest area in the radar chart, demonstrating that both ATPL and Co-Learnability are necessary. The ATPL only variant (brown) outperforms ACCEL (blue) on 9 out of the 12 held-out mazes, confirming that incorporating transition prediction error into the regret approximation enhances the agent’s generalization ability. In contrast, the Co-Learnability only variant (teal) falls below the ACCEL baseline, indicating that Co-Learnability is effective only when paired with our regret approximation. Detailed numerical results are provided in Appendix E. Additional ablations on BipedalWalker and sensitivity analyses for $\alpha$ and $\beta$ appear in Appendix B.
5 Conclusion
In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition prediction error term for regret approximation, and (ii) a lightweight Co-Learnability metric that captures cross-task transfer effects. By integrating these into the standard generator–replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
Empirically, we demonstrated on two procedurally generated domains (MiniGrid navigation and BipedalWalker) that TRACED outperforms baselines (DR, PLR, ADD, ACCEL), including sota CENIE, using only half the training updates. We showed superior zero-shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes. Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp-up, while co-learnability yields gains when paired with our regret estimates.
Looking forward, Co-Learnability offers a simple, computationally light mechanism for capturing inter-task influences and could be further refined via more sophisticated causal estimators or learned models. More broadly, any RL setting that relies on regret-approximation stands to benefit from incorporating transition-prediction error, providing an easy-to-implement boost in sample efficiency. We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret-guided exploration in open-ended learning environments.
Appendix A Proof
A.1 Proof of Proposition 1
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proof. Since $\mathcal{T}_{t+1}=\Lambda$,
$\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}\;=\;\frac{1}{|\Lambda|}\sum_{j\in\Lambda}Y_{j}=\mathrm{CoLearnability}(i,t)$
A.2 Proof of Theorem 2
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
Proof. Let
$q_{s}\;=\;\Pr\bigl{[}\text{choose }i\text{ at time }t_{0}+s\bigr{]}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
so that $1-q_{s}$ is the probability $i$ is not chosen at step $s$.
Since for every $j\in\Lambda$ we have $\mathrm{TaskPriority}(j,t_{0}+s)\leq M_{0})$, it follows that
$\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)\;\leq\;|\Lambda|\;M_{0}.$
Hence
$q_{s}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j}\mathrm{TaskPriority}(j,t_{0}+s)}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{|\Lambda|\;M_{0}}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0})}{|\Lambda|\;M_{0}}=:\;c,$ (8)
For any nonnegative integer $k$,
$\Pr[T>k]$ $=\Pr\bigl{(}\text{task }i\text{ is not chosen in any of steps }0,1,\dots,k-1\bigr{)}$
$=\prod_{s=0}^{k-1}\bigl{(}1-q_{s}\bigr{)}\;\leq\;\prod_{s=0}^{k-1}\bigl{(}1-c\bigr{)}=\bigl{(}1-c\bigr{)}^{k}.$
where the inequality uses (8).
$\mathbb{E}[T]=\sum_{k=0}^{\infty}\Pr[T>k]\;\leq\;\sum_{k=0}^{\infty}(1-c)^{k}\;=\;\frac{1}{c}\;=\;\frac{|\Lambda|\;M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}.$
This completes the proof.
Appendix B Additional Experiments
B.1 Ablation Study of Scaling Factor
In this ablation, we keep every setting in Appendix H fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight). Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains as shown in Table 2. For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance. The default TRACED weights show the best performance compared to fixed projection baselines. This result shows that an effective balance between ATPL and Co-Learnability is needed.
Table 2: MiniGrid Ablation Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks, comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$) over 5 seeds. For TRACED, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Each method is evaluated after 10k PPO updates. Bold indicates the best result per task; underline indicates the second-best.
Environment $\alpha$=0.0 $\alpha$=100.0 $\beta$=0.0 $\beta$=100.0 TRACED
16Rooms 0.81 $\pm$ 0.1 0.06 $\pm$ 0.04 0.73 $\pm$ 0.09 0.74 $\pm$ 0.17 0.79 $\pm$ 0.19
16Rooms2 0.94 $\pm$ 0.05 0.0 $\pm$ 0.0 0.28 $\pm$ 0.2 0.59 $\pm$ 0.15 0.72 $\pm$ 0.17
SimpleCrossing 0.84 $\pm$ 0.04 0.43 $\pm$ 0.08 0.86 $\pm$ 0.04 0.82 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.45 $\pm$ 0.05 0.23 $\pm$ 0.04 0.41 $\pm$ 0.05 0.52 $\pm$ 0.03 0.47 $\pm$ 0.02
SmallCorridor 0.42 $\pm$ 0.24 0.12 $\pm$ 0.06 0.62 $\pm$ 0.21 0.51 $\pm$ 0.12 0.49 $\pm$ 0.17
LargeCorridor 0.44 $\pm$ 0.26 0.04 $\pm$ 0.02 0.54 $\pm$ 0.16 0.56 $\pm$ 0.18 0.5 $\pm$ 0.14
Labyrinth 0.5 $\pm$ 0.29 0.0 $\pm$ 0.0 0.49 $\pm$ 0.25 0.19 $\pm$ 0.18 1.0 $\pm$ 0.0
Labyrinth2 0.61 $\pm$ 0.2 0.0 $\pm$ 0.0 0.2 $\pm$ 0.17 0.2 $\pm$ 0.2 0.98 $\pm$ 0.01
Maze 0.74 $\pm$ 0.21 0.0 $\pm$ 0.0 0.18 $\pm$ 0.07 0.39 $\pm$ 0.24 0.59 $\pm$ 0.17
Maze2 0.56 $\pm$ 0.26 0.0 $\pm$ 0.0 0.51 $\pm$ 0.21 0.14 $\pm$ 0.12 0.42 $\pm$ 0.19
Maze3 0.51 $\pm$ 0.27 0.0 $\pm$ 0.0 0.59 $\pm$ 0.22 0.51 $\pm$ 0.21 0.86 $\pm$ 0.07
PerfectMaze(M) 0.4 $\pm$ 0.16 0.01 $\pm$ 0.02 0.4 $\pm$ 0.06 0.34 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.6 $\pm$ 0.1 0.07 $\pm$ 0.01 0.5 $\pm$ 0.09 0.46 $\pm$ 0.05 0.7 $\pm$ 0.04
B.2 Ablation Study of ATPL and Co-Learnability
Table 3 breaks down zero-shot solved rates on 6 held-out BW tasks for TRACED and its ablations at 10k updates. TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one. The ATPL only variant (TRACED – CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED – ATPL) only reaches the second place in two tasks (Hardcore, Roughness). Removing both components (the ACCEL baseline) ranks second on one (Stairs). These results confirm that both transition prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
Table 3: BipedalWalker Ablation Results. Mean $\pm$ SE of zero-shot test returns on 6 held-out BipedalWalker tasks, comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$) over 5 seeds. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
Basic 281.65 $\pm$ 5.25 282.61 $\pm$ 4.64 286.75 $\pm$ 5.01 293.67 $\pm$ 3.56
Hardcore 37.59 $\pm$ 15.0 53.77 $\pm$ 20.8 48.35 $\pm$ 23.74 86.83 $\pm$ 17.96
Stairs -38.71 $\pm$ 10.54 -43.99 $\pm$ 12.99 -34.1 $\pm$ 8.44 -29.0 $\pm$ 10.4
PitGap -65.07 $\pm$ 7.57 -72.92 $\pm$ 13.95 -5.2 $\pm$ 66.89 -39.26 $\pm$ 11.42
Stump -79.18 $\pm$ 5.45 -38.06 $\pm$ 39.88 -19.89 $\pm$ 64.11 34.16 $\pm$ 54.58
Roughness 161.72 $\pm$ 28.36 191.75 $\pm$ 25.07 191.68 $\pm$ 4.99 193.29 $\pm$ 21.6
Mean 49.67 $\pm$ 11.24 62.19 $\pm$ 11.21 77.93 $\pm$ 17.58 89.95 $\pm$ 12.95
B.3 Hyperparameter Sensitivity Analysis
Table 4 reports the impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED on 12 held-out MiniGrid mazes, each evaluated after 10k PPO updates. Four of the six $(\alpha,\beta)$ settings outperform the 10k update baselines (Appendix E), and even the weakest setting matches the performance of PLR at 10k. This demonstrates that TRACED delivers consistently strong zero-shot generalization across a wide range of hyperparameters, without the need for extensive tuning.
Table 4: Effect of Hyperparameter Weights $(\alpha,\beta)$ on MiniGrid Generalization. Mean $\pm$ SE of zero-shot solved rates across 12 held-out MiniGrid environments, evaluated under various $(\alpha,\beta)$ settings. Except for the final column, TRACED with $(\alpha=1.0,\beta=1.0)$, averaged over five seeds, all other entries report single-seed measurements. Bold indicates the best result per task; underline indicates the second-best.
Environment (0.75, 1.0) (1.25, 1.0) (1.0, 0.75) (1.0, 1.25) (1.25, 0.75) TRACED (1.0, 1.0)
16Rooms 0.26 0.90 1.00 0.96 1.00 $0.79\pm 0.19$
16Rooms2 0.00 0.05 0.08 0.78 0.10 $\underline{0.72}\pm 0.17$
SimpleCrossing 0.82 0.84 0.85 0.71 0.78 $\textbf{0.89}\pm 0.01$
FourRooms 0.68 0.36 0.52 0.39 0.51 $0.47\pm 0.02$
SmallCorridor 0.55 0.00 0.01 0.00 0.01 $\underline{0.49}\pm 0.17$
LargeCorridor 0.94 0.02 0.01 0.00 0.01 $\underline{0.50}\pm 0.14$
Labyrinth 0.98 1.00 1.00 1.00 1.00 $\textbf{1.00}\pm 0.00$
Labyrinth2 1.00 0.73 0.76 1.00 0.95 $\underline{0.98}\pm 0.01$
Maze 0.76 0.16 0.39 0.00 1.00 $0.59\pm 0.17$
Maze2 0.07 0.00 0.94 1.00 0.98 $0.42\pm 0.19$
Maze3 0.92 1.00 0.10 0.89 0.93 $0.86\pm 0.07$
PerfectMaze(M) 0.56 0.38 0.55 0.55 0.85 $\underline{0.66}\pm 0.11$
Mean 0.63 0.45 0.52 0.61 0.67 $\textbf{0.70}\pm 0.04$
Table 5 reports zero-shot test returns on six held-out BipedalWalker terrains after 10k PPO updates, under five different $(\alpha,\beta)$ settings. Notably, three of those configurations outperform all 10k-update baselines (Appendix E). Even the mean return across all five $(\alpha,\beta)$ choices ($\approx 64.9$) exceeds every baseline. These results demonstrate that TRACED delivers consistently strong zero-shot generalization across a wide range of weight settings, without requiring extensive hyperparameter tuning.
Table 5: Effect of Hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker Generalization. Mean $\pm$ SE of zero-shot test returns on six held-out BipedalWalker terrains, for various $(\alpha,\beta)$ settings. Only the final column (TRACED with $(\alpha=1.5,\beta=0.6)$) is averaged over five seeds; all other entries are single-seed runs. Bold indicates the best result per terrain; underline indicates the second-best.
Environment (1.25, 0.6) (1.75, 0.6) (1.5, 0.45) (1.5, 0.75) TRACED (1.5, 0.6)
Basic 282.06 291.04 281.32 279.37 293.67$\pm$3.56
Hardcore –7.60 115.87 29.55 50.10 86.83$\pm$17.96
Stairs –52.78 –21.94 –42.37 –66.02 –29.0$\pm$10.4
PitGap 12.96 –52.72 –52.21 –122.54 –39.26$\pm$11.42
Stump –100.95 –40.11 –74.83 –73.93 34.16$\pm$54.58
Roughness 148.22 187.43 244.57 185.83 193.29$\pm$21.6
Mean 46.99 79.93 64.34 42.13 89.95$\pm$12.95
Appendix C Efficiency Analysis
C.1 Wall-Clock Training Time in BipedalWalker
Table 6: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration (hours $\pm$ SE over five runs) for each algorithm at 10k and 20k PPO updates.
Domain TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
BipedalWalker 35.64 $\pm$ 0.53 33.26 $\pm$ 1.17 70.22 $\pm$ 5.9 21.13 $\pm$ 0.11 41.58 $\pm$ 0.24 35.49 $\pm$ 1.07 71.01 $\pm$ 2.17 18.38 $\pm$ 0.1 37.03 $\pm$ 0.3
Table 6 reports wall-clock training times on the BipedalWalker domain. At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 6% more than ACCEL’s $33.26\pm 1.17$h, yet delivers a 22% increase in mean solved rate (see Section 4).
C.2 Sample Complexity: Environment Steps
Table 7: Environment Steps. Total environment interactions (millions) for each method per given number of student PPO updates.
Environment PPO Updates PLR ADD ACCEL TRACED
MiniGrid 10k 82M 41M 93M 93M
MiniGrid 20k 165M 82M 185M
BipedalWalker 10k 165M 80M 174M 174M
BipedalWalker 20k 329M 160M 347M
Table 7 shows total environment steps required to achieve a fixed number of PPO updates. TRACED matches ACCEL’s sample complexity exactly, confirming that its superior generalization stems from improved curriculum design rather than additional data.
Appendix D Visualizing Curriculum Dynamics
(a) 2.5k, Low ATPL, Low Co-Learnability
(b) 2.5k, Low ATPL, High Co-Learnability
(c) 2.5k, High ATPL, Low Co-Learnability
(d) 2.5k, High ATPL, High Co-Learnability
(e) 5k, Low ATPL, Low Co-Learnability
(f) 5k, Low ATPL, High Co-Learnability
(g) 5k, High ATPL, Low Co-Learnability
(h) 5k, High ATPL, High Co-Learnability
Figure 12: Representative terrains from BipedalWalker selected by ATPL and co-learnability at two training stages. The top two rows show terrains after 2.5k PPO updates (first row: low ATPL; second row: high ATPL), and the bottom two rows after 5k updates (third row: low ATPL; fourth row: high ATPL). Within each row, the left column displays low co-learnability examples and the right column high co-learnability examples.
Figure 12 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at two training milestones (2.5k and 5k PPO updates). ATPL alone captures the raw challenge level: low-ATPL terrains (first and third rows) remain relatively smooth, while high-ATPL terrains (second and fourth rows) feature larger gaps and steeper bumps. As training proceeds from 2.5k to 5k updates, the overall terrains become systematically harder, demonstrating TRACED’s ability to ramp up difficulty in lockstep with the agent’s improving skills. Co-Learnability then refines this progression by filtering out extremes: low Co-Learnability (left column) tends to generate either trivial or overwhelmingly difficult levels that can stall learning, whereas high Co-Learnability (right column) favors intermediate-difficulty terrains that transfer more effectively across tasks. Together, these visualizations confirm that ATPL drives a steadily increasing complexity trajectory, while Co-Learnability smooths it to sustain robust, transferable learning throughout the curriculum.
Appendix E Numerical Results
E.1 Minigrid Environment
Table 8: MiniGrid Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks for DR, PLR, ADD, ACCEL, and TRACED, evaluated after 10k and 20k PPO updates over 10 seeds. Bold indicates the best result per task; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
16Rooms 10k 0.4 $\pm$ 0.12 0.81 $\pm$ 0.06 0.72 $\pm$ 0.1 0.76 $\pm$ 0.12 0.53 $\pm$ 0.14
20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
16Rooms2 10k 0.1 $\pm$ 0.09 0.42 $\pm$ 0.11 0.14 $\pm$ 0.05 0.48 $\pm$ 0.13 0.58 $\pm$ 0.13
20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
SimpleCrossing 10k 0.58 $\pm$ 0.05 0.72 $\pm$ 0.03 0.63 $\pm$ 0.03 0.81 $\pm$ 0.03 0.85 $\pm$ 0.03
20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
FourRooms 10k 0.32 $\pm$ 0.04 0.47 $\pm$ 0.01 0.39 $\pm$ 0.03 0.46 $\pm$ 0.02 0.46 $\pm$ 0.02
20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
SmallCorridor 10k 0.3 $\pm$ 0.07 0.52 $\pm$ 0.13 0.47 $\pm$ 0.1 0.37 $\pm$ 0.14 0.49 $\pm$ 0.13
20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
LargeCorridor 10k 0.36 $\pm$ 0.13 0.33 $\pm$ 0.13 0.35 $\pm$ 0.09 0.44 $\pm$ 0.15 0.43 $\pm$ 0.11
20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
Labyrinth 10k 0.1 $\pm$ 0.1 0.5 $\pm$ 0.15 0.22 $\pm$ 0.13 0.6 $\pm$ 0.15 0.93 $\pm$ 0.03
20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
Labyrinth2 10k 0.1 $\pm$ 0.1 0.28 $\pm$ 0.13 0.0 $\pm$ 0.0 0.52 $\pm$ 0.14 0.96 $\pm$ 0.03
20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
Maze2 10k 0.19 $\pm$ 0.12 0.43 $\pm$ 0.14 0.09 $\pm$ 0.08 0.44 $\pm$ 0.15 0.6 $\pm$ 0.1
20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
Maze3 10k 0.27 $\pm$ 0.13 0.57 $\pm$ 0.13 0.2 $\pm$ 0.1 0.78 $\pm$ 0.13 0.81 $\pm$ 0.06
20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
PerfectMaze(M) 10k 0.06 $\pm$ 0.02 0.24 $\pm$ 0.05 0.17 $\pm$ 0.08 0.41 $\pm$ 0.05 0.64 $\pm$ 0.07
20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
Mean 10k 0.23 $\pm$ 0.05 0.45 $\pm$ 0.04 0.28 $\pm$ 0.04 0.54 $\pm$ 0.03 0.66 $\pm$ 0.03
20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
Table 9: MiniGrid Ablation Results. Mean $\pm$ SE of zero-shot solved rates on 12 held-out MiniGrid tasks, comparing ACCEL (10k), TRACED – ATPL (10k), TRACED – CL (10k), and TRACED (10k) over 5 seeds. For TRACED and ACCEL, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
Table 10: PerfectMaze Zero-Shot Solved Rates. Mean $\pm$ SE of zero-shot success rates on PerfectMazeLarge and PerfectMazeXL after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, and TRACED over 10 seeds. Bold indicates the best result per environment; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
PerfectMaze(Large) 10k 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01 0.04 $\pm$ 0.04 0.09 $\pm$ 0.03 0.27 $\pm$ 0.07
20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
PerfectMaze(XL) 10k 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.02 0.01 $\pm$ 0.01 0.1 $\pm$ 0.04
20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
Table 8 reports per-task zero-shot solved rates on the 12 held-out MiniGrid mazes, following [21]. We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (at both 10k and 20k). TRACED at 10k achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best rate in one maze (Maze2). Averaged across all 12 mazes, TRACED attains a mean solved rate of 0.66, outperforming every other baseline configuration.
Table 9 presents zero-shot solved rates on 12 held-out MiniGrid mazes for TRACED and its ablations at 10k updates. Full TRACED attains the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the other six, with a mean solved rate of 0.70, outperforming every variant. These results confirm that both transition prediction error and Co-Learnability are critical to TRACED’s superior, consistent performance across diverse tasks.
Table 10 compares zero-shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large-scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL. After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third-highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k. These results underscore TRACED’s ability to scale to extreme maze sizes.
E.2 BipedalWalker Environment
Table 11 compares TRACED’s zero-shot returns against DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains. At 10k updates, TRACED attains the highest mean return in three environments (Basic, Hardcore, and Stump) and achieves an overall mean of 89.95 $\pm$ 31.72, outperforming all baselines.
Table 11: BipedalWalker Zero-Shot Test Returns. Mean $\pm$ SE of episode returns on six held-out BipedalWalkerHardcore terrains after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, ACCEL-CENIE, and TRACED over 5 seeds. Bold indicates the best result per terrain; underline indicates the second-best. ACCEL-CENIE is estimated by applying the original CENIE paper’s ACCEL-to-ACCEL-CENIE performance ratio [20] to our ACCEL implementation.
Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
Basic 10k 112.56 $\pm$ 62.99 131.63 $\pm$ 56.58 119.43 $\pm$ 63.72 281.65 $\pm$ 5.25 273.56 $\pm$ 1.5 293.67 $\pm$ 3.56
20k 68.17 $\pm$ 44.71 97.67 $\pm$ 61.73 63.64 $\pm$ 62.71 281.85 $\pm$ 3.72 275.04 $\pm$ 3.19
Hardcore 10k -19.67 $\pm$ 7.35 -17.36 $\pm$ 5.99 3.83 $\pm$ 9.07 37.59 $\pm$ 15.0 66.83 $\pm$ 17.48 86.83 $\pm$ 17.96
20k -16.98 $\pm$ 5.8 -23.7 $\pm$ 2.05 10.48 $\pm$ 20.12 59.23 $\pm$ 25.5 84.09 $\pm$ 34.32
Stairs 10k -17.6 $\pm$ 7.21 -7.23 $\pm$ 7.16 -5.87 $\pm$ 9.07 -38.71 $\pm$ 10.54 -30.4 $\pm$ 8.35 -29.0 $\pm$ 10.4
20k -9.11 $\pm$ 7.67 -10.2 $\pm$ 4.5 -0.66 $\pm$ 9.2 -46.34 $\pm$ 5.78 -36.39 $\pm$ 11.48
PitGap 10k -34.39 $\pm$ 15.66 -46.21 $\pm$ 15.69 -17.8 $\pm$ 13.82 -65.07 $\pm$ 7.57 -39.81 $\pm$ 18.75 -39.26 $\pm$ 11.42
20k -26.03 $\pm$ 10.98 -48.72 $\pm$ 12.23 -7.65 $\pm$ 7.52 -64.89 $\pm$ 18.93 -45.92 $\pm$ 33.9
Stump 10k -24.99 $\pm$ 6.76 -27.74 $\pm$ 6.33 -26.15 $\pm$ 8.49 -79.18 $\pm$ 5.45 -60.05 $\pm$ 10.96 34.16 $\pm$ 54.58
20k -25.52 $\pm$ 12.39 -23.61 $\pm$ 2.86 3.26 $\pm$ 19.1 -67.18 $\pm$ 15.56 -46.34 $\pm$ 176.21
Roughness 10k -0.66 $\pm$ 12.01 3.55 $\pm$ 9.91 20.61 $\pm$ 13.78 161.72 $\pm$ 28.36 174.92 $\pm$ 2.25 193.29 $\pm$ 21.6
20k 0.19 $\pm$ 8.18 -2.73 $\pm$ 9.51 18.71 $\pm$ 21.59 213.48 $\pm$ 7.69 224.4 $\pm$ 8.63
Mean 10k 2.54 $\pm$ 15.44 6.1 $\pm$ 10.69 14.4 $\pm$ 14.05 49.67 $\pm$ 11.24 64.18 89.95 $\pm$ 12.95
20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
Table 12 compares zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED. A trajectory is considered “solved” if its return exceeds 230. At 10k updates, TRACED achieves the top solved rate in five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36—substantially higher than ACCEL’s 0.29 and the other baselines. No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
Table 12: BipedalWalker Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates, defined as the fraction of episodes exceeding a return of 230, on 6 held-out BipedalWalkerHardcore terrains (Basic, Hardcore, Stairs, PitGap, Stump, Roughness) after 10k and 20k PPO updates. Results compare DR, PLR, ADD, ACCEL, and TRACED over 5 seeds. Bold indicates the best per terrain; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
Basic 10k 0.35 $\pm$ 0.2 0.43 $\pm$ 0.21 0.34 $\pm$ 0.21 0.98 $\pm$ 0.02 1.00 $\pm$ 0.0
20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
Hardcore 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.01 $\pm$ 0.01 0.15 $\pm$ 0.03 0.28 $\pm$ 0.05
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
Roughness 10k 0.01 $\pm$ 0.01 0.0 $\pm$ 0.01 0.01 $\pm$ 0.01 0.56 $\pm$ 0.1 0.65 $\pm$ 0.08
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
Mean 10k 0.06 $\pm$ 0.04 0.07 $\pm$ 0.04 0.06 $\pm$ 0.04 0.29 $\pm$ 0.02 0.36 $\pm$ 0.03
20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03
Appendix F Implementation Details
F.1 Minigrid Environment
SixteenRooms
SixteenRooms2
SmallCorridor
LargeCorridor
SimpleCrossing
FourRooms
Labyrinth
Labyrinth2
Maze
Maze2
Maze3
PerfectMaze
Figure 13: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid [16], and PerfectMaze from REPAIRED [6].
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid [16]. Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three—turn left, turn right, and move forward—are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work [4, 6]. Figure 13 displays all evaluation tasks.
F.2 BipedalWalker Environment
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym [23]. At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of $-100$ is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
We evaluate agents in the BipedalWalker-v3 environment, which features relatively smooth terrain, as well as in the more demanding BipedalWalkerHardcore-v3, which introduces a variety of challenging obstacles. The BipedalWalker-v3 focuses on basic locomotion and balance without significant terrain disturbances. Terrain in this environment does not have obstacles such as stairs, pit gaps, and stumps, and only gentle slopes are generated by adding noise in the range of approximately -1 to 1. In contrast, BipedalWalkerHardcore-v3 includes complex terrain elements such as stairs, pit gaps, stumps, and roughness, which require precise control and adaptive strategies. These obstacles are procedurally generated with randomized parameters. The stair height is randomly set to either $+1$ (ascending) or $-1$ (descending), with the number of steps and step width randomly sampled between 3–5 steps and 4–5 units, respectively. The pit width is determined by sampling a single integer between 3–5 units. Both the height and width of each stump are sampled as a single random integer between 1–3 units, ensuring the stumps are always square-shaped. Terrain roughness noise range is the same as BipedalWalker-v3.
To further probe the agent’s generalization capabilities, we include four targeted evaluation environments that isolate specific terrain challenges: Stairs, PitGap, Stump, and Roughness. Stairs specifies a fixed stair height of $2$ units, with $5$ steps and a step width between $4$ and $5$. PitGap sets the pit width to exactly $5$ units. Stump defines the stump height as $2$ units while allowing the stump width to vary between $1$ and $2$ units. Finally, Roughness generates terrain by adding noise $5$. Figure 14 displays all evaluation tasks.
(a) BipedalWalker-v3
(b) BipedalWalkerHardcore-v3
(c) Stairs
(d) PitGap
(e) Stump
(f) Roughness
Figure 14: Evaluation Terrains in BipedalWalker. Six held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
F.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128-dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video-prediction tasks under partial observability [25, 26].
For walking in challenging terrain (BipedalWalker), the model instead concatenates the raw state vector $s_{t}$ with $a_{t}$ and feeds this into an LSTM (hidden size 128, 1 layer); the recurrent output is then linearly mapped to the predicted next state $\hat{s}_{t+1}$.
Its instantaneous error is defined by
$\ell_{t}=\begin{cases}\|\,\hat{s}_{t+1}-s_{t+1}\|_{1},&\text{in MiniGrid environments},\\ \|\hat{s}_{t+1}-s_{t+1}\|_{2}^{2},&\text{in BipedalWalker}.\end{cases}$
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction [27], while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains [28].
The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
For partially observable navigation, the input consists of a 3-channel $5\times 5$ local image observation and a directional scalar with dimension 4 indicating the agent’s current orientation. The image is processed by a convolutional encoder with a filter size of 16, a kernel size of 3, and Rectified Linear Unit (ReLU) activations, followed by flattening. The directional scalar is passed through a fully connected layer of size 5, and the resulting embedding is concatenated with the flattened image features. The combined feature vector is fed separately to the actor and critic heads, each consisting of fully connected layers of size [32, 32]. The actor outputs a categorical distribution over discrete actions, while the critic produces a scalar value estimate $V(s)$.
For walking in a challenging terrain (BipedalWalker), the input is a flat state vector of dimension 24. This input is processed through a Multilayer Perceptron (MLP) with two hidden layers of 64 dimensions, followed by actor and critic heads. The actor head outputs the mean and standard deviation parameters of a diagonal Gaussian distribution over a continuous 4-dimensional action space, and the critic head outputs a scalar state value $V(s)$. The PPO learning process optimizes a clipped surrogate policy loss with clip parameter $\epsilon$ = 0.2 combined with a value function loss.
Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
$\widehat{\mathrm{Regret}}(\tau)\;=\;\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau).$
If the buffer is full, we evict the lowest-priority task before appending. Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret. We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer. Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq. (7)) before proceeding to the next timestep.
Algorithm 1 UED Workflow with Transition-Prediction Loss and Co-Learnability
1:Given: policy $\pi_{\phi}$, level generator $\mathcal{G}$, buffer capacity $K$, replay probability $p_{\mathrm{replay}}$, batch size $B$, mutation count $N_{\text{mutate}}$, scaling factors $\alpha,\beta$
2:Initialize Task Buffer $\Lambda\leftarrow\emptyset$, timestep $t\leftarrow 0$
3:while policy has not converged do
4:     Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
5:     if $d_{t}=0$ $\triangleright$ Exploration phase
6:         Sample new task $\theta\sim\mathcal{G}$
7:         Collect trajectory $\tau$ by executing $\pi_{\phi}$ on $\theta$ (no gradient)
8:         Compute regret estimate
$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)+\alpha\,\mathrm{ATPL}(\tau)$
9:         Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest-priority)
10:     $\triangleright$ Replay + Mutation phase
11:         Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\mathrm{TaskPriority}(\cdot,t)$
12:         for $k=1\ldots B$ do
13:              Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
14:              Update policy $\phi$ using rewards from $\tau_{k}$
15:              Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
16:         end for
17:         Select the $N$ tasks $\{\theta_{k^{\prime}}\}$ with smallest $\widehat{\mathrm{Regret}}(\tau_{k^{\prime}})$
18:         for each selected $\theta_{k^{\prime}}$ do
19:              $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
20:              Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
21:              Compute $\widehat{\mathrm{Regret}}(\tilde{\tau})$
22:              Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
23:         end for
24:     end if
25:     Buffer update:
26:     for each task $i\in\Lambda$ updated at $t$ do
27:         Recompute $\mathrm{TaskDifficulty}(i,t)$
28:         Recompute $\mathrm{CoLearnability}(i,t-1)$
29:         Update $\mathrm{TaskPriority}(i,t)$ via Eq. (7)
30:     end for
31:     $t\leftarrow t+1$
32:end while
Appendix H Hyperparameters
Table 13 contains the hyperparameters we use. Hyperparameter values were adopted directly from the DCD repository [29]. We perform hyperparameter tuning in MiniGrid and Bipedal environments, focusing on two key components, $\alpha$ and $\beta$. We explored values in the ranges $\alpha\in[0.5,1.0,1.5]$, $\beta\in[0.6,0.8,1.0]$.
Table 13: Hyperparameters used for training each method in each environment.
HyperParameter MiniGrid BipedalWalker
PPO
$\gamma$ 0.995 0.99
$\lambda_{\text{GAE}}$ 0.95 0.9
PPO rollout length 256 2048
PPO epochs 5 5
PPO minibatches per epoch 1 32
PPO clip range 0.2 0.2
PPO number of workers 16 4
Adam learning rate 1e-4 3e-4
Adam $\epsilon$ 1e-5 1e-5
PPO max gradient norm 0.5 0.5
PPO value clipping True False
Return normalization False True
Value loss coefficient 0.5 0.5
Student entropy coefficient 0.0 1e-3
Generator entropy coefficient 0.0 1e-2
TRACED
ATPL weight, $\alpha$ 1.0 1.5
Co-Learnability weight, $\beta$ 1.0 0.6
ACCEL
Replay rate, $p_{\text{replay}}$ 0.8 0.9
Buffer size, $K$ 4000 1000
Number of edits 5 3
Batch Size, $B$ 4 4
Number of mutated tasks, $N_{\text{mutate}}$ 4 4
Temperature 0.3 0.1
Staleness coefficient, $\rho$ 0.3 0.5
Appendix I Level Evolution
I.1 Visualization of Level Evolution in Minigrid
The visualization of how the Minigrid environment evolves as the number of blocks increases is shown in Figure 15. Each step of the evolutionary process produces an edited level that has a high learning efficiency.
Figure 15: Visualization of a single level’s evolving progression in the MiniGrid environment. Starting from top-left, ending bottom-right. This progress is automatically designed by our TRACED algorithm.
I.2 Evolution of Levels in BipedalWalker
Figure 16 shows the complexity metric results trained by 3 methods, TRACED (Ours), ACCEL + ATPL, ACCEL. Starting with plain terrain (near zero point), all 3 methods guide levels with more complex terrain. With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL. This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
(a) Stump height low
(b) Stump height high
(c) Stair height low
(d) Stair height high
(e) Stair height step
(f) Pit gap low
(g) Pit gap high
(h) Roughness
Figure 16: Emergent BipedalWalker terrain complexity metrics. Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue). The result is averaged over 5 random seeds. This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
I.3 Visualization of Level Evolution in BipedalWalker
Figure 17: Visualization of the level evolving progression in the BipedalWalker environment. Starting from top-left, ending bottom-right. in this example, starting with plain terrain, the pits are created and their number increases, then the roughness increases, then stairs and stumps are created and their number, width, and height increase. This progress is automatically designed by our TRACED algorithm.
Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases. As we can see, the process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
Appendix J Agent Trajectory Visualizations Across Environments
J.1 MiniGrid
Figure 18 shows a representative rollout in the PerfectMaze environment.
Figure 18: Agent trajectory visualization in the PerfectMaze evaluation task. TRACED enables efficient path planning and robust generalization to challenging maze environments, as demonstrated by the agent’s successful navigation to the goal (green).
J.2 BipedalWalker
Figure 19 illustrates a zero-shot rollout in BipedalWalkerHardcore. The agent dynamically modulates joint torques to ascend and descend steps, leap over gaps, step onto stumps, and maintain balance on uneven ground, confirming TRACED’s effectiveness in teaching transferrable locomotion skills.
Figure 19: Agent trajectory visualization in the BipedalWalkerHardcore. TRACED agent successfully overcomes various combined obstacles, including stairs, pit gaps, stumps, and roughness, demonstrating robust generalization to complex environments.
Appendix K Difference with Former Studies
K.1 Difference with ACCEL
Our approach builds on ACCEL’s evolutionary, regret-based curriculum design, refining both its regret estimator and its sampling criterion. Like ACCEL, we maintain a level buffer and record each task’s difficulty. However, instead of relying solely on Positive Value Loss (PVL) to approximate regret, we decompose regret into value error and transition-prediction error and incorporate both components into our difficulty measure.
Furthermore, prior UED methods sample levels independently based solely on their individual difficulty, ignoring cross-task transfer effects. We address this with our Co-Learnability metric, which quantifies how training on one task influences regret on others by measuring the average change in their difficulty after replay. By integrating Co-Learnability into our priority score, we can prioritize tasks that not only challenge the agent but also deliver broad transfer benefits across the entire task space.
K.2 Difference with CENIE
CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL. It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state–action space. This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
However, CENIE’s diversity-driven objective differs from ours. TRACED’s primary goal is to approximate task difficulty more faithfully, by decomposing regret into value and transition-prediction errors, and to capture how training on one task transfers to others via our Co-Learnability metric. While CENIE accounts for per-task novelty and regret, it ignores cross-task transfer effects. Moreover, TRACED is agnostic to the choice of level generator or editor, meaning our method could be combined with CENIE’s novelty scoring to yield a curriculum that balances both transfer-aware difficulty and environmental diversity.
Appendix L Related Work
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments [24, 30, 4, 19]. PAIRED [4] trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Mediratta et al. [9] introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr [31] employs variational autoencoder–based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
Another line of work is based on Prioritized Level Replay (PLR) [5], which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  [6] make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL [7] extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods [5, 6, 7] approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
Moving beyond regret-based approaches, CENIE [20] integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR [32] introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state–action trajectories. MBeDED [33] introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO [34] extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON [10] applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE [35] suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA [36] is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL [37] aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task [38]. Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress [39]. Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process [40, 41]. WAKER [42] leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
Appendix M More Details On Baseline Algorithms
In Unsupervised Environment Design (UED), Domain Randomization (DR) constructs the curriculum by uniformly sampling tasks from the environment parameter space $\Theta$. Formally, DR samples tasks according to $\theta\sim p(\Theta)$ where $p(\Theta)$ represents a uniform distribution over the task parameter space. The agent is trained on these randomly sampled tasks.
Robust PLR (PLR[6] was introduced as an extension of the original PLR [5]. While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
Adversarially Compounding Complexity by Editing Levels (ACCEL) [7] actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
Coverage-based Evaluation of Novelty In Environment (CENIE) [20] introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state–action pairs and the learned distribution of past state–action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) [19] frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase [29], and assessed ADD with the ADD codebase [19]. CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
Appendix N Long-term Analysis on TRACED
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 20: TRACED Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 21: ACCEL Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
TRACED reaches near-peak solved rates very quickly. However, its performance subsequently oscillates rather than remaining perfectly stable (Figure 20). Crucially, this same post-peak fluctuation also appears under the ACCEL curriculum (Figure 21), where the agent simply arrives more slowly at its maximum but then exhibits the identical up-and-down pattern. In other words, these oscillations reflect an inherent tendency of the RL policy to suffer occasional catastrophic failures in held-out levels, rather than a pathology introduced by our task-prioritization strategy. Addressing and damping these fluctuations in the student agent’s performance is an important direction for future work.
Appendix O Limitations
As with any research, our approach presents several limitations that highlight opportunities for future investigation:
  • Co-Learnability In this paper, we proposed Co-Learnability, simple yet effective method. Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
  • Sampling levels We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods. However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
  • Heuristic weighting We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty. While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
  • Level mutation We adopted the same approach as ACCEL which employs random mutations. Our approach focused on improving task priority rather than advancing environment parameter mutation mechanisms. The current random mutation process creates a clear opportunity for future work to advance these methods using evolutionary algorithms such as CMA-ES [43], so the Editor can offer the agent more informative tasks.
  • Extension to other RL algorithms Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC [44] and TD3 [45] feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 [46] build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
  • Extend to other UED methods We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR [6] and CENIE [20]) and conduct comprehensive experiments to assess its performance in these settings.
  • Evaluation in other domains Our experimental validation has been conducted in MiniGrid [16, 4] and BipedalWalker [23, 7]. Evaluating TRACED in other domains such as Car Racing [23, 6], physical robotics applications [47], and reasoning task ARC [48] would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
Appendix P Experimental Setup and Reproducibility
All experiments were conducted using Python 3.8.20 on Ubuntu 22.04.4 LTS. The hardware setup included an AMD EPYC 7543 32-core processor, 80GB of RAM, and an NVIDIA A100 GPU. Experiments were run on a shared server with 8 GPUs, with each experiment using a single GPU. We implemented all experiments in Python 3.8.20, using PyTorch(v1.9.0+cu111). The environment stack was built using MiniGrid (v1.0.1), Gym (v0.15.7).

### v3

@@
 Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge.
-One promising solution is Unsupervised Environment Design (UED), a co-evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum.
-Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value-function loss.
-Building on these approaches, we introduce the transition prediction error as an additional term in our regret approximation.
+One promising solution is Unsupervised Environment Design (UED), a co‑evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum.
+Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value‑function loss.
+Building on these approaches, we introduce the transition-prediction error as an additional term in our regret approximation.
 To capture how training on one task affects performance on others, we further propose a lightweight metric called Co-Learnability.
-By combining these two measures, we present Transition-aware Regret Approximation with Co-learnability for Environment Design (TRACED).
-Empirical evaluations show that TRACED yields curricula that improve zero-shot generalization across multiple benchmarks while requiring up to 2$\times$ fewer environment interactions than strong baselines.
-Ablation studies confirm that the transition prediction error drives rapid complexity ramp-up and that co-learnability delivers additional gains when paired with the transition prediction error.
+By combining these two measures, we present Transition‑aware Regret Approximation with Co‑learnability for Environment Design (TRACED).
+Empirical evaluations show that TRACED produces curricula that improve zero-shot generalization over strong baselines across multiple benchmarks.
+Ablation studies confirm that the transition-prediction error drives rapid complexity ramp‑up and that Co‑Learnability delivers additional gains when paired with the transition-prediction error.
 These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED.
+https://geonwoo.me/traced/
 1 Introduction
-Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics [1].
+Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (Sutton et al., 1998).
 Ideally, we want agents that generalize robustly to a broad range of unseen environments.
-However, hand-crafting a training distribution that captures all real-world variability is intractable, and agents often overfit even large training sets, performing poorly out-of-distribution [2, 3].
-Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent [4].
-A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design [4, 5, 6, 7, 8, 9, 10].
+However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (Kirk et al., 2023; Korkmaz, 2024).
+Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (Dennis et al., 2020).
+A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022; Azad et al., 2023b; Mediratta et al., 2023; Erlebach & Cook, 2024).
 Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains.
-Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) [11].
-In this paper, we refine regret estimation by explicitly decomposing it into (i) a value error, captured by PVL, and (ii) a transition prediction error, estimated via a transition dynamics model.
-We further introduce a metric called Co-Learnability to quantify how training on one task benefits others.
-For instance, consider three 100-word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy [12, 13, 14], reflecting high Co-Learnability; in contrast, Japanese is typologically distant [15], so gains from Japanese may transfer unefficiently to English, reflecting low Co-Learnability.
-We present a lightweight estimator of Co-Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
-Figure 1: Task Priority Landscape. Task in the top-right quadrant (high difficulty and high co-learnability) are scheduled with the highest priority in the curriculum.
-As illustrated in Figure 1, these two ingredients, refined regret and Co-Learnability, are fused into a unified Task Priority score via a rank-based transform (cf.
-[7]), which governs both new task generation and prioritized replay.
-We call our algorithm TRACED (Transition-aware Regret Approximation with Co-learnability for Environment Design).
-We evaluate TRACED on procedurally generated MiniGrid (MG) [16] and BipedalWalker (BW) [17].
-In MG, we compare against Domain Randomization (DR) [18], PLR [5], ADD [19], and ACCEL [7], the strongest of these baselines.
-In BW, we additionally include the state-of-the-art (sota) CENIE [20].
-On 12 held-out MG mazes, TRACED achieves a mean solved rate of 66% after 10k updates, outperforming ACCEL’s 60% after 20k updates.
-In the extreme maze variants PerfectMaze, TRACED reaches 27% success on PerfectMazeLarge and 10% on PerfectMazeXL, representing 35-40% relative gains over ACCEL with only half the training time.
-In BW, TRACED achieves an 18% improvement over CENIE while halving the wall-clock time.
-Ablation studies reveal that the transition-prediction error term drives rapid complexity ramp-up, and that Co-Learnability provides additional gains when combined with our regret term.
-Collectively, these results pave a new path toward sample-efficient curriculum design in UED by combining refined regret approximation with explicit modeling of task relationships.
+Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (Rutherford et al., 2024).
+In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics.
+This combined signal provides a more faithful approximation of regret and a sharper basis for curriculum design.
+We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others.
+For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (Ramírez et al., 2013; Costa et al., 2000; Lemhöfer & Dijkstra, 2004), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick & and, 2005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability.
+We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
+Figure 1: Task Priority Landscape. Task with high difficulty and high Co‑Learnability are scheduled with the highest priority in the curriculum.
+We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1).
+We evaluate TRACED on procedurally generated MiniGrid (MG) (Chevalier-Boisvert et al., 2023) and BipedalWalker (BW) (Romac et al., 2021).
+In MG, we compare against Domain Randomization (DR) (Jakobi, 1997), PLR (Jiang et al., 2021b), ADD (Chung et al., 2024), and ACCEL (Parker-Holder et al., 2022) (the strongest baseline).
+In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Teoh et al., 2024).
+TRACED surpasses all baselines in mean solved rate across 12 MiniGrid mazes, BipedalWalker, and even the extreme PerfectMaze variants.
+Ablation studies show that the transition-prediction error term accelerates curriculum ramp-up, while Co-Learnability provides additional gains when combined with our regret term.
+An analysis of curriculum evolution indicates that TRACED progressively increases task difficulty from easy to challenging.
+Taken together, these results chart a path toward more sample-efficient UED by coupling refined regret approximation with explicit modeling of task relationships.
 We provide the full implementation of TRACED to facilitate the reproduction of our main results at https://github.com/Cho-Geonwoo/TRACED.
@@
 We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al.
-[4].
+(2020).
 A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space.
@@
 Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
-At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory-conditioned policy $\pi\bigl{(}a_{t}|o_{0},a_{0},\dots,o_{t}\bigr{)}$.
+At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory‑conditioned policy $\pi\bigl(a_{t}|o_{0},a_{0},\dots,o_{t}\bigr)$.
 For a fixed level $\theta$, the utility of policy $\pi$ is the expected discounted return $U_{\theta}(\pi)=\mathbb{E}\!\Bigl[\sum_{t=0}^{T}\gamma^{t}\,r_{t}\Bigr],r_{t}=\mathcal{R}(s_{t},a_{t})$ with the expectation taken over both the stochastic dynamics and the policy’s choices.
@@
 2.2 Unsupervised Environment Design (UED)
-UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ [4].
-Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning [5, 7, 20].
+UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (Dennis et al., 2020).
+Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (Jiang et al., 2021b; Parker-Holder et al., 2022; Teoh et al., 2024).
 The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
 Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics.
-For example, PLR [6] evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
-PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)-based Temporal Difference (TD) errors over an episode.
+For example, PLR (Jiang et al., 2021a) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
+PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode.
 PVL for an episode $\tau$ of length $T$ is defined as
 $\mathrm{PVL}(\tau)=\frac{1}{T}\sum_{t=0}^{T}\max({\sum_{k=t}^{T}(\gamma\lambda)^{\,k-t}\,\delta_{k}},0\Bigr)$ (1)
-where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one-step TD-error at timestep $t$.
+where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one‐step TD‐error at timestep $t$.
 MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target.
-Other criteria such as policy entropy, one-step TD error, GAE, policy min-margin, and policy least-confidence have also been evaluated in curriculum learning contexts [5].
-3 Method
-At each step, curriculum design aims to select the tasks that will most effectively drive the agent’s learning progress.
-Our approach enhances UED by combining (1) a refined regret approximation integrating both value and transition errors and (2) a simple yet principled measure of task co-learnability.
-We compute a unified Task Priority score from these two quantities, which then governs both new task generation and replay sampling in the UED loop.
+Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (Jiang et al., 2021b).
+3 TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design
+TRACED improves UED via (i) a regret approximation combining value and transition-prediction errors and (ii) a Co-Learnability measure.
+These yield a unified Task Priority score that governs new-task generation and replay sampling.
+We first present the motivating regret decomposition, then the curriculum derived from the approximated regret.
 3.1 Regret Approximation via Transition Prediction Loss
-Since regret quantifies the difficulty an agent experiences on a task, improving regret approximation can lead to more accurate task difficulty estimation and thus more effective curricula.
-To address this, we decompose one-step regret at state–action pair $(s,a)$ into two interpretable terms:
-$\mathrm{Regret}(s,a)$ $=Q^{*}(s,a)-Q(s,a)$ (2)
-$=\gamma\Bigl{(}\mathbb{E}_{s_{1}\sim P(\cdot\mid s,a)}\bigl{[}Q^{*}(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}(\cdot\mid s,a)}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}\Bigr{)}$
-$=\gamma\Bigl{(}\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}\,Q^{*}(s_{1},a^{\prime})-Q(s_{1},a^{\prime})\,\bigr{]}}_{\text{(i) Value Error}}+\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}Q(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}}_{\text{(ii) Transition Prediction Error}}\Bigr{)}$
-Term (i), the value error, measures the expected discrepancy between the true optimal action-value $Q^{*}$ and the agent’s learned $Q$.
-Term (ii), the transition prediction error, quantifies the bias introduced by replacing the true transition kernel $P$ with an empirical approximation $\hat{P}$, where $\hat{P}(s^{\prime}\!\mid s,a)$ is the conditional distribution estimated from the agent’s collected experience $\{(s,a,s^{\prime})\}$.
-Accurately estimating both terms is crucial: term (i) reflects the agent’s value-approximation quality, whereas term (ii) captures how model-mismatch in the dynamics affects long-term returns.
-Because $Q^{*}$ and $P$ are unknown, we approximate (i) via the Positive Value Loss (PVL) and (ii) via a learned transition dynamics estimator $f_{\phi}$. Concretely:
+Since regret quantifies the difficulty an agent experiences on a task, improving its approximation can yield more accurate difficulty estimates and, in turn, more effective curricula.
+We approximate the one-step regret at a state-action pair $(s,a)$ via the following decomposition:
+$\mathrm{Regret}(s,a)$ $=V^{*}(s)-Q^{\pi}(s,a)$ (2)
+$=V^{*}(s)-\hat{V}^{*}(s)+\hat{V}^{*}(s)-Q^{\pi}(s,a)$
+$=\underbrace{V^{*}(s)-\hat{V}^{*}(s)}_{\text{(i) Value estimation error}}+\underbrace{r(s,a^{*})-r(s,a)}_{\text{(ii) Reward gap}}$
+$\quad+\gamma\,\underbrace{\Bigl(\mathbb{E}_{s^{\prime\prime}\sim\hat{P}(\cdot\mid s,a^{*})}\!\bigl[\hat{V}^{*}(s^{\prime\prime})\bigr]-\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\!\bigl[V^{\pi}(s^{\prime})\bigr]\Bigr)}_{\text{(iii) Future value gap}}\,$,
+Notation.
+A hat (e.g., $\hat{V}^{*}(s)$) indicates an empirical/learned estimator of the optimal value; $Q^{*},V^{*}$ are the optimal value functions; $Q^{\pi},V^{\pi}$ are the value functions under policy $\pi$; $P$ is the true transition kernel and $\hat{P}$ is a learned transition model; $\gamma$ is the discount factor.
+Eq.
+2 makes clear that Positive Value Loss (PVL), which evaluates only the accuracy of the empirical value estimator and thus corresponds to term (i), is insufficient as a proxy for regret.
+The future-value gap in term (iii) is influenced not only by value-function error but also by mismatch between the learned dynamics $\hat{P}$ and the true dynamics $P$.
+Motivated by this, we augment PVL with a transition-prediction error term, explicitly accounting for model-environment dynamics mismatch when approximating regret.
 Definition 1 (Average Transition Prediction Loss).
-Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one-step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ (details in Appendix F.3).
+Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one‐step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ .
+Detailed design choices for the transition model are provided in Appendix H.3.
 Over an episode $\tau=(s_{0},a_{0},\dots,s_{T})$, define
@@
 $\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau),$ (4)
-where $\alpha>0$ balances value versus transition terms (see Appendix B).
-By explicitly incorporating both components, our regret approximation more faithfully reflects task difficulty and yields more effective curricula.
-After computing the approximated regret $\widehat{\mathrm{Regret}}(\tau)$ for each sampled task instance, we append this value to the Task Difficulty Buffer (TDB).
-Consequently, the number of entries in the TDB corresponding to task $i$ exactly equals the number of times task $i$ has been sampled (see Figure 2 for an overview).
+where $\alpha>0$ balances value versus transition terms.
+By explicitly incorporating both components, our regret approximation more faithfully captures task difficulty and, in turn, yields more effective curricula.
+For each sampled task instance $\tau$, we compute $\widehat{\mathrm{Regret}}(\tau)$ and append it to the task difficulty buffer (TDB), which is subsequently used to drive curriculum updates (Figure 2).
 Figure 2: Task Difficulty Calculation Workflow.
@@
 The agent interacts with sampled tasks to collect episode trajectories, which are stored in the Rollouts Buffer (RB).
-For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition-Prediction Loss (ATPL).
+For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition‑Prediction Loss (ATPL).
 Their sum produces the updated task difficulty (approximated regret), which is appended to the TDB to refresh each sampled task’s stored difficulty.
 3.2 Task Priority Construction
-A central challenge in UED is to decide which tasks to present to the agent at each step.
+A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes et al., 2024).
 Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others.
-To this end, we introduce two complementary quantities, Task Difficulty and Co-Learnability, and combine them into a single Task Priority score.
+To this end, we introduce two complementary quantities, Task Difficulty and Co‑Learnability, and combine them into a single Task Priority score.
+In the following sections, time $t$ indexes the teacher’s curriculum-selection cycles (not environment time).
 Definition 2 (Task Difficulty).
-Let $s_{i}(t)\;=\;\max\{\,s\leq t:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $t$ when task $i$ was sampled (and its approximated regret stored).
-Then we define
-$\mathrm{TaskDifficulty}(i,t)=\begin{cases}\mathrm{TDB}\bigl{(}i,\,s_{i}(t)\bigr{)},&\text{if }s_{i}(t)\text{ is finite},\\[4.0pt] 0,&\text{if }i\text{ has never been sampled before }t.\end{cases}$ (5)
-This ensures that $\mathrm{TaskDifficulty}(i,t)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
-Definition 3 (Co-Learnability).
+Let $s_{i}(k)\;=\;\max\{\,s\leq k:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $k$ when task $i$ was sampled, so that its approximated regret was stored at time $s_{i}(k)$.
+We define the task difficulty of task $i$ at time $k$ as
+[…]
+This ensures that $\mathrm{TaskDifficulty}(i,k)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
+Definition 3 (Co‑Learnability).
 Beyond difficulty, we wish to capture how training on one task accelerates progress on others.
-Let $\mathcal{T}_{t+1}$ be the set of tasks replayed at time $t+1$.
-We define
-$\mathrm{CoLearnability}(i,t)=\frac{1}{|\mathcal{T}_{t+1}|}\sum_{j\in\mathcal{T}_{t+1}}\bigl{[}\mathrm{TaskDifficulty}(j,t)-\mathrm{TaskDifficulty}(j,t+1)\bigr{]}$ (6)
-i.e.,  the average reduction in difficulty of the replayed tasks induced by sampling $i$. Positive co-learnability indicates that visiting $i$ yields broad transfer benefits.
-Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
-$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
-the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
-$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
-Proposition 1 confirms that our co-learnability definition precisely measures the mean one-step change in task difficulties, conditional on having replayed task i, when the next sampled task set covers the entire buffer.
+Let $\mathcal{T}_{k+1}$ be the set of tasks replayed at time $k+1$.
+We define the Co-Learnability (CL) of task i at time k as
+$\mathrm{CoLearnability}_{i}(k)=\frac{1}{|\mathcal{T}_{k+1}|}\sum_{j\in\mathcal{T}_{k+1}}\bigl[\mathrm{TaskDifficulty}(j,k)-\mathrm{TaskDifficulty}(j,k+1)\bigr]$ (6)
+which measures the average reduction in the difficulty of replayed tasks when task $i$ is selected at time $k$.
+In principle, the true marginal contribution could be computed via Shapley values; however, due to computational constraints we approximate task $i$’s effect on reducing other tasks’ difficulty using this surrogate.
+A positive Co-Learnability value indicates that visiting $i$ yields transfer benefits.
 Definition 4 (Task Priority).
-We combine difficulty and Co-Learnability into a scalar score, then pass it through a rank function, which replaces raw values by their relative ordering, to obtain robust sampling weights as in ACCEL [7]:
+We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (Parker-Holder et al., 2022):
 $\mathrm{TaskPriority}(i,t)=\mathrm{Rank}\Bigl(\mathrm{TaskDifficulty}(i,t)+\beta\,\mathrm{CoLearnability}(i,t)\Bigr)$ (7)
 where $\beta>0$ trades off challenge versus transfer.
-We sample tasks for replay proportionally to their computed priority scores, so that higher-priority tasks are more likely to be selected at each step.
-By ranking, we prevent outliers from dominating the curriculum and ensure a smooth sampling distribution.
-Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
-$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
-[…]
-Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
-$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
-This guarantees that no task remains “stale” indefinitely: its expected waiting time scales inversely with its initial priority.
-While the boundedness and non-increasing requirements in Theorem 3.2 may initially seem restrictive, they follow naturally from the agent’s continual learning process.
-Indeed, Figure 10 shows an overall upward trend in returns during training, indicating that the agent’s performance steadily improves and, consequently, task priority scores tend to decrease over time.
+At each step, we sample tasks inversely to their priority (e.g., $p(i\mid t)\propto 1/\mathrm{TaskPriority}(i,t)$), so lower ranks (higher priority) are selected more often.
+The rank transform prevents outliers from dominating the sampling distribution.
 3.3 Overall UED Workflow
 Figure 3: Method Workflow Overview.
-The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer in proportion to their priority scores.
+The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer based on their priority scores.
 (2) Priority Update: we recompute each level’s priority based on our task priority definition (Eq.
 7).
-(3) Task Mutation: the lowest-priority levels are mutated into new variants and reinserted into the buffer.
-The overall UED algorithm follows the same loop as ACCEL [7], with the sole modification that task scoring uses our task priority definition (Eq.
-7) in place of positive value loss alone.
-We maintain a policy $\pi_{\phi}$, a level generator G, a Task Buffer $\Lambda$, and the sampling distribution $P_{D}$ over the decision variable $d\in\{0,1\}$, where $d=0$ indicates generator sampling and $d=1$ indicates buffer replay.
-At each iteration $t$, we draw $d_{t}\sim P_{D}$ and proceed as follows.
-If $d_{t}=0$, we sample a fresh task parameter $\theta$ from the generator and run the current policy to collect a trajectory $\tau$ on $\theta$.
-We then compute the approximated regret score $\widehat{\mathrm{Regret}}(\tau)$ using Eq.
-4.
-If $\widehat{\mathrm{Regret}}(\tau)$ exceeds a predefined threshold, $\theta$ is appended to $\Lambda$.
-If instead $d_{t}=1$, we sample a batch of $B$ tasks $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ according to their TaskPriority distribution (Eq.
-7).
-For each $\theta_{k}$, we execute the policy $\pi_{\phi}$ to collect trajectory $\tau_{k}$, perform the standard RL update on $\phi$, and compute its approximated regret $\widehat{\mathrm{Regret}}(\tau_{k})$.
-We then select the $N_{\text{mutate}}$ tasks with the lowest regret scores and feed each through the task editor to obtain a mutated variant $\tilde{\theta}$.
-Finally, we replace the corresponding entries in $\Lambda$ with these $\tilde{\theta}$, ensuring that easier tasks are systematically evolved into more challenging ones.
-Following each update to the Task Difficulty Buffer, we recompute the $\mathrm{TaskDifficulty}(i,t)$ for every modified task $i$ and update the $\mathrm{CoLearnability}(j,t-1)$ for all previously sampled tasks $j$.
-These refreshed scores determine the priorities for the next replay phase.
-The algorithm thus alternates between sampling new tasks and performing prioritized replay until the policy performance converges.
-The full procedure is detailed in Algorithm 1.
+(3) Task Mutation: the lowest‑priority levels are mutated into new variants and reinserted into the buffer.
+The overall UED algorithm follows the ACCEL loop (Parker-Holder et al., 2022), with the sole change that task scoring uses our Task Priority (Eq.
+7) in place of Positive Value Loss (PVL) alone (Figure 3).
+At each curriculum update time $t$, with probability $d$ we sample a new level uniformly at random, and with probability 1d1-d we sample a level from the replay buffer according to its Task Priority score.
+Early in training, when the buffer is empty, we sample levels uniformly at random for a warm-up period.
+The warm-up length matches ACCEL’s setting.
+After training the agent on the selected level, we approximate its regret using the agent’s value-related loss and the transition-prediction loss (for Task Difficulty), and append the result to the task difficulty buffer.
+Co-Learnability is updated following Eq.
+6 after the next curriculum step, once the newly visited levels have had their difficulties updated.
+The updated Task Difficulty and Co-Learnability define Task Priority via Eq.
+7, which then governs both new-task sampling and prioritized replay at future times.
+The algorithm alternates between sampling new tasks and prioritized replay until policy performance converges.
+The full procedure is given in Algorithm 1.
 4 Experiments
@@
 In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD.
-In BW only, we additionally include the recent sota method CENIE [20].
-Detailed descriptions of all baselines can be found in Appendix M.
-For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero-shot test performance on held-out levels.
-To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library [21].
+In BW only, we additionally include the recent sota method CENIE (Teoh et al., 2024).
+We exclude SFL (Rutherford et al., 2024) because, on MiniGrid, it does not outperform PLR (Rutherford et al., 2024).
+Detailed descriptions of all baselines can be found in Appendix K.
+For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels.
+To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal et al., 2021).
 The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible.
 Accordingly, higher IQM and lower optimality gap values indicate better performance.
-All methods use Proximal Policy Optimization (PPO) [22] as the student agent.
-Following prior work [7, 20], we plot the performance against the number of PPO updates; the corresponding environment interactions per update are reported in Appendix C.
-For more details on hyperparameters and network structures, refer to Appendix F.
+All methods use Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the student agent.
+Following prior work (Parker-Holder et al., 2022; Teoh et al., 2024), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F.
+Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16.
+Consequently, some baseline scores may differ from those in the ACCEL paper.
+However, Appendix A.4 shows that TRACED has no worker-specific tuning and that reducing workers does not render the comparison unfair.
+We report TRACED at 10k updates (rather than 20k) because single-seed, long-horizon runs (45k PPO updates) on MiniGrid reveal post-convergence oscillations in both TRACED and ACCEL, and reporting earlier avoids this confound (Appendix B).
+Additional hyperparameters and architectures are provided in Appendix H.
 SixteenRooms
@@
 LargeCorridor
-(a) Maze
+(a) Minigrid
 Stairs
@@
 (b) BipedalWalker
-Figure 5: Held-out Evaluation Environments.
-(a) Example held-out mazes for zero-shot evaluation: 4 tasks are shown (see Appendix F.1 for all 12 task definitions).
-(b) Example held-out BipedalWalker terrains for zero-shot evaluation: 2 tasks are shown (see Appendix F.2 for all 6 task definitions).
+Figure 5: Held‑out Evaluation Environments.
+(a) Example held‑out MiniGrid mazes for zero‑shot evaluation: 4 tasks are shown (see Appendix H.1 for all 12 task definitions).
+(b) Example held‑out BipedalWalker terrains for zero‑shot evaluation: 2 tasks are shown (see Appendix H.2 for all 6 task definitions).
 4.1 Partially Observable Navigation
-We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid [16], as in prior UED work.
+We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (Chevalier-Boisvert et al., 2023), as in prior UED work.
 Four representatives are shown in Figure 5.
-The agent observes a 147-dimensional pixel observation and is trained for up to 20k PPO updates.
-Figure 6: Zero-Shot Transfer Performance in MiniGrid.
-Aggregated solved rates on held-out MiniGrid mazes after 10k and 20k PPO updates.
-TRACED at 10k updates (red) matches or outperforms DR (orange), $\text{PLR}^{\perp}$ (green), ADD (gray), and ACCEL (blue) at 20k updates.
-Averaged over 10 random seeds.
-As shown in Figure 6, our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines.
-In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and co-learnability scoring.
-The mean performance mirrors this ordering, showing that even lower-performing seeds see substantial gains.
-Furthermore, TRACED at 10k produces the smallest optimality gap, closing the distance to an ideal target success rate faster than any competitor; in contrast, ACCEL requires only 20k updates to halve its gap.
-Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain for each algorithm at 10k and 20k PPO updates.
+The agent observes a 147‑dimensional pixel observation and is trained for up to 20k PPO updates.
+Figure 6: Zero‑Shot Transfer Performance in MiniGrid. Aggregated solved rates on held‑out MiniGrid mazes after 10k and 20k PPO updates. TRACED at 10k updates outperforms baselines at 20k updates.
+Our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines (Figure 6).
+In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and Co‑Learnability scoring.
+Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain.
 TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
 13.78 $\pm$ 0.36 12.94 $\pm$ 0.66 26.58 $\pm$ 0.76 22.48 $\pm$ 0.27 45.08 $\pm$ 0.29 14.87 $\pm$ 0.62 31.83 $\pm$ 1.36 5.82 $\pm$ 0.12 12.41 $\pm$ 0.18
-Compared to ACCEL, TRACED halves the wall-clock training time while maintaining equivalent or better transfer performance (Table 1).
+Compared to ACCEL, TRACED halves the wall‑clock training time while maintaining equivalent or better transfer performance (Table 1).
 Even relative to ACCEL at the same 10k update budget, TRACED incurs a 6% computational overhead yet delivers a 22% relative increase in mean solved rate.
-Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero-shot transfer across diverse MiniGrid mazes.
-Detailed per-task zero-shot results are provided in Appendix E, and the number of environment interactions per PPO update for each method is in Appendix C.
+Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero‑shot transfer across diverse MiniGrid mazes.
+Detailed per‑task zero‑shot results are provided in Appendix P, and the number of environment interactions per PPO update for each method is in Appendix F.
 (a) Shortest Path Length
@@
 Shortest path length and number of blocks both grow faster under TRACED (pink) than ACCEL (blue).
-This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
-Averaged over 10 random seeds with shaded 95% confidence intervals.
-To analyze emergent environment complexity, we track two structural metrics for each generated maze: the length of the shortest solution path and the number of obstacles (blocks), as plotted in Figure 7.
-Both metrics rise much more rapidly under TRACED (red) than under ACCEL (blue), showing that our priority scoring more effectively distinguishes easy versus challenging tasks to produce a steadily escalating curriculum.
-Notably, the ACCEL + ATPL (brown) also drives complexity upward far faster than ACCEL.
-It closely parallels TRACED’s gains, demonstrating that the transition-prediction component alone contributes greatly to complexity ramp-up.
+This faster ramp‑up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
+To analyze emergent environment complexity, we track two structural metrics for each generated maze: (i) the length of the shortest solution path and (ii) the number of obstacles (Figure 7).
+Curves are averaged over 10 seeds with shaded 95% confidence intervals.
+Both metrics increase substantially faster under TRACED than under ACCEL, indicating that our priority scoring more effectively separates easy from challenging tasks and yields a steadily escalating curriculum.
+Notably, ACCEL+ATPL also drives complexity upward far faster than ACCEL alone and closely tracks TRACED, demonstrating that the transition-prediction component on its own contributes strongly to complexity ramp-up.
 (a) PerfectMazeLarge
@@
 Figure 8: PerfectMaze Evaluation.
-(a), (b) Two held-out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero-shot testing.
-(c) Zero-shot solved rates after 10k and 20k PPO updates (mean $\pm$ std over 10 random seeds).
+(a), (b) Two held‑out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero‑shot testing.
+(c) Zero‑shot solved rates.
 TRACED achieves the highest 10k performance on PerfectMazeLarge and closely matches the best 20k performance on PerfectMazeXL.
-To stress-test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks.
+To stress‐test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks.
 PerfectMazeLarge consists of $51\times 51$ grids with a maximum episode length exceeding 5k steps, while PerfectMazeXL scales this to $100\times 100$ grids.
 Figure 8 shows representative levels from each variant.
-We evaluate zero-shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total).
-As depicted in Figure 8, on PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$.
+We evaluate zero‐shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total).
+On PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$ (Figure 8).
 On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes.
-Detailed per-baseline results for both benchmarks are provided in Appendix E.
+Detailed per-baseline results for both benchmarks are provided in Appendix P.
 4.2 Walking in Challenging Terrain
-We further validate our curriculum in the continuous-control BipedalWalkerHardcore environment from OpenAI Gym [23], as modified by Wang et al.
-(2019) [24].
+We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman et al., 2016), as modified by Wang et al.
+(2019).
 This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty.
@@
 Figure 5 illustrates two representative terrains with varying stair heights and surface roughness.
-Figure 9: Aggregate zero-shot performance on BipedalWalker terrains.
+We evaluate zero‑shot transfer on six held‑out test terrains over 100 episodes each, averaged over five random seeds.
+Figure 9: Aggregate zero‑shot performance on BipedalWalker terrains.
 All scores are normalized by the maximum return of 300 in the BipedalWalker domain.
 TRACED at 10k updates (pink) matches or exceeds ACCEL-CENIE at 20k updates (purple) across all metrics.
-Averaged over 5 random seeds.
 (a) Basic
@@
 (f) Roughness
-Figure 10: Return progression on BipedalWalker terrains.
-TRACED (pink) consistently outperforms DR (orange), PLR (green), ADD (gray), and ACCEL (blue).
-Averaged over 10 random seeds with shaded 95% confidence intervals.
-Our student agent is trained for 10k PPO updates, while baselines run up to 20k updates.
-We evaluate zero-shot transfer on six held-out test terrains over 100 episodes each.
-Figure 9 reveals consistent gains across all four aggregate metrics, median, IQM, mean, and optimality gap, on the six held-out BipedalWalker terrains.
-Specifically, after only 10k updates, TRACED outperforms all baselines at 20k updates.
-Although its variance is slightly higher, TRACED’s 10k-update curve matches or exceeds ACCEL-CENIE at 20k, demonstrating both superior sample efficiency and robust generalization.
-Moreover, as in MG, TRACED reaches the highest performance in half the wall-clock time on BW relative to ACCEL (see Appendix C).
-These results confirm that our curriculum design remains effective not only in discrete grid worlds but also in complex continuous domains.
-Refer to Appendix E for detailed results broken down by each terrain.
-As shown in Figure 10, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains.
-In the first 5k updates, TRACED trails ADD, DR, and PLR because it builds upon ACCEL’s seed curriculum, mutating the easiest environments first, before increasing difficulty.
-After 5k updates, however, TRACED continues to climb steadily, while ADD, DR, and PLR plateau in a stale regime.
-This sustained upward trajectory demonstrates that TRACED effectively escalates task complexity in tandem with the agent’s improving capabilities, yielding superior long-term performance.
+Figure 10: Return progression on BipedalWalker terrains. TRACED consistently outperforms baselines.
+TRACED delivers consistent gains across all four aggregate metrics, median, interquartile mean (IQM), mean, and optimality gap, on the six held-out BipedalWalker terrains (Figure 9).
+After only 10k updates, it already outperforms all baselines evaluated at 20k.
+As in MiniGrid, TRACED also reaches its peak performance in roughly half the wall-clock time of ACCEL on BipedalWalker (Appendix F).
+Moreover, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains throughout training (Figure 10), further underscoring the effectiveness of its curriculum design.
+Detailed numerical results are provided in Appendix P.
 4.3 Ablation Study
-Figure 11: Both ATPL and CL are important design choices.
-Showing per-task zero-shot solved rates for 4 variants: full TRACED (pink), ATPL only (brown), Co-Learnability only (teal), and ACCEL (blue).
-Averaged over 5 random seeds.
-To isolate the contributions, we compare TRACED at 10k updates against three ablations (Figure 11): ATPL only (TRACED – CL), Co-Learnability only (TRACED – ATPL), and the ACCEL baseline (TRACED – ATPL – CL).
-TRACED (pink) encloses the largest area in the radar chart, demonstrating that both ATPL and Co-Learnability are necessary.
-The ATPL only variant (brown) outperforms ACCEL (blue) on 9 out of the 12 held-out mazes, confirming that incorporating transition prediction error into the regret approximation enhances the agent’s generalization ability.
-In contrast, the Co-Learnability only variant (teal) falls below the ACCEL baseline, indicating that Co-Learnability is effective only when paired with our regret approximation.
-Detailed numerical results are provided in Appendix E.
-Additional ablations on BipedalWalker and sensitivity analyses for $\alpha$ and $\beta$ appear in Appendix B.
+Figure 11: Ablation Study on TRACED. Both ATPL and CL are important design choices.
+To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds.
+Results are reported following Agarwal et al.
+(2021).
+On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance.
+The same trend holds on BipedalWalker (Appendix A).
+Ablations on the scaling factors $\alpha$ and $\beta$ further show that both elements play significant roles in TRACED and that carefully balancing ATPL and CL can yield additional gains (Appendix A.1).
+Detailed numerical results are provided in Appendix P.
+4.4 Analysis on Curriculum Progression
+Table 2: Proportions (%) of Easy, Moderate, and Challenging levels in the level buffer during PPO updates.
+Method Difficulty 0k 5k 10k 15k 20k
+ACCEL Easy $100$ $100$ $79.2$ $26.7$ $9.2$
+Moderate 0 0 $20.8$ $73.3$ $90.8$
+Challenging 0 0 0 0 0
+TRACED Easy $100$ $72$ $24.5$ $19.5$ $11.2$
+Moderate 0 $25.7$ $60.9$ $68$ $77.3$
+Challenging 0 $2.3$ $14.6$ $12.5$ $11.5$
+To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters.
+Following Teoh et al.
+(2024), we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5).
+Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging.
+Table 2 reports the proportion of each difficulty in the buffer at different PPO update steps.
+TRACED progressively shifts mass from Easy to Moderate and Challenging, introducing nontrivial proportions of Challenging levels by 10k (14.6%) and maintaining them thereafter (11.5% at 20k).
+In contrast, ACCEL never surfaces Challenging levels even at 20k (0% throughout), with the buffer dominated by Moderate levels by 20k (90.8%).
+This steady escalation under TRACED aligns with its faster convergence and stronger performance already at 10k updates.
+The respective contributions of ATPL and Co-Learnability (CL) to curriculum progression are analyzed in Appendix C, and a visualization of level evolution is provided in Appendix D.
 5 Conclusion
-In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition prediction error term for regret approximation, and (ii) a lightweight Co-Learnability metric that captures cross-task transfer effects.
-By integrating these into the standard generator–replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
+In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition-prediction error term for regret approximation, and (ii) a lightweight Co‑Learnability metric that captures cross‐task transfer effects.
+By integrating these into the standard generator-replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
 Empirically, we demonstrated on two procedurally generated domains (MiniGrid navigation and BipedalWalker) that TRACED outperforms baselines (DR, PLR, ADD, ACCEL), including sota CENIE, using only half the training updates.
-We showed superior zero-shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes.
-Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp-up, while co-learnability yields gains when paired with our regret estimates.
-Looking forward, Co-Learnability offers a simple, computationally light mechanism for capturing inter-task influences and could be further refined via more sophisticated causal estimators or learned models.
-More broadly, any RL setting that relies on regret-approximation stands to benefit from incorporating transition-prediction error, providing an easy-to-implement boost in sample efficiency.
-We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret-guided exploration in open-ended learning environments.
-Appendix A Proof
-A.1 Proof of Proposition 1
-Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
-$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
-the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
-$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
-Proof. Since $\mathcal{T}_{t+1}=\Lambda$,
-$\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}\;=\;\frac{1}{|\Lambda|}\sum_{j\in\Lambda}Y_{j}=\mathrm{CoLearnability}(i,t)$
-A.2 Proof of Theorem 2
-Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
-$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
-[…]
-Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
-$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
-Proof. Let
-$q_{s}\;=\;\Pr\bigl{[}\text{choose }i\text{ at time }t_{0}+s\bigr{]}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
-so that $1-q_{s}$ is the probability $i$ is not chosen at step $s$.
-Since for every $j\in\Lambda$ we have $\mathrm{TaskPriority}(j,t_{0}+s)\leq M_{0})$, it follows that
-$\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)\;\leq\;|\Lambda|\;M_{0}.$
-Hence
-$q_{s}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j}\mathrm{TaskPriority}(j,t_{0}+s)}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{|\Lambda|\;M_{0}}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0})}{|\Lambda|\;M_{0}}=:\;c,$ (8)
-For any nonnegative integer $k$,
-$\Pr[T>k]$ $=\Pr\bigl{(}\text{task }i\text{ is not chosen in any of steps }0,1,\dots,k-1\bigr{)}$
-$=\prod_{s=0}^{k-1}\bigl{(}1-q_{s}\bigr{)}\;\leq\;\prod_{s=0}^{k-1}\bigl{(}1-c\bigr{)}=\bigl{(}1-c\bigr{)}^{k}.$
-where the inequality uses (8).
-$\mathbb{E}[T]=\sum_{k=0}^{\infty}\Pr[T>k]\;\leq\;\sum_{k=0}^{\infty}(1-c)^{k}\;=\;\frac{1}{c}\;=\;\frac{|\Lambda|\;M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}.$
-This completes the proof.
-Appendix B Additional Experiments
-B.1 Ablation Study of Scaling Factor
-In this ablation, we keep every setting in Appendix H fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight).
-Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains as shown in Table 2.
-For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance.
-The default TRACED weights show the best performance compared to fixed projection baselines.
-This result shows that an effective balance between ATPL and Co-Learnability is needed.
-Table 2: MiniGrid Ablation Solved Rates.
-Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks, comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$) over 5 seeds.
-For TRACED, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation.
+We showed superior zero‑shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes.
+Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp‑up, while Co‑Learnability yields gains when paired with our regret estimates.
+Looking forward, Co‑Learnability offers a simple, computationally light mechanism for capturing inter‐task influences and could be further refined via more sophisticated causal estimators or learned models.
+More broadly, any RL setting that relies on regret‑approximation stands to benefit from incorporating transition‐prediction error, providing an easy‐to‐implement boost in sample efficiency.
+We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret‐guided exploration in open‑ended learning environments.
+Appendix A Additional Experiments
+A.1 Ablation Study on the Scaling Factor
+In this ablation, we keep every setting in Appendix O fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight).
+Results are averaged over five random seeds and reported as mean $\pm$ standard error on 12 held-out MiniGrid tasks.
 Each method is evaluated after 10k PPO updates.
-Bold indicates the best result per task; underline indicates the second-best.
+Table 3: Ablation study on the scaling factor.
+Comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$).
+Bold indicates the best; underline indicates the second-best.
 Environment $\alpha$=0.0 $\alpha$=100.0 $\beta$=0.0 $\beta$=100.0 TRACED
@@
 Mean 0.6 $\pm$ 0.1 0.07 $\pm$ 0.01 0.5 $\pm$ 0.09 0.46 $\pm$ 0.05 0.7 $\pm$ 0.04
-B.2 Ablation Study of ATPL and Co-Learnability
-Table 3 breaks down zero-shot solved rates on 6 held-out BW tasks for TRACED and its ablations at 10k updates.
-TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one.
-The ATPL only variant (TRACED – CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED – ATPL) only reaches the second place in two tasks (Hardcore, Roughness).
-Removing both components (the ACCEL baseline) ranks second on one (Stairs).
-These results confirm that both transition prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
-Table 3: BipedalWalker Ablation Results.
-Mean $\pm$ SE of zero-shot test returns on 6 held-out BipedalWalker tasks, comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$) over 5 seeds.
+Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains (Table 3).
+For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance.
+The default TRACED weights show the best performance compared to fixed projection baselines.
+This result shows that an effective balance between ATPL and Co-Learnability is needed.
+A.2 Ablation Study on ATPL and Co-Learnability
+In this ablation, we evaluate TRACED by removing each of its two components in turn.
+Table 4 reports zero-shot solve rates on six held-out BipedalWalker tasks after 10k PPO updates.
+Results are averaged over five random seeds and reported as mean $\pm$ standard error.
+Table 4: Ablation study on ATPL and Co-Learnability.
+Comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$).
 Bold indicates the best result per task; underline indicates the second-best.
@@
 Mean 49.67 $\pm$ 11.24 62.19 $\pm$ 11.21 77.93 $\pm$ 17.58 89.95 $\pm$ 12.95
-B.3 Hyperparameter Sensitivity Analysis
-Table 4 reports the impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED on 12 held-out MiniGrid mazes, each evaluated after 10k PPO updates.
-Four of the six $(\alpha,\beta)$ settings outperform the 10k update baselines (Appendix E), and even the weakest setting matches the performance of PLR at 10k.
-This demonstrates that TRACED delivers consistently strong zero-shot generalization across a wide range of hyperparameters, without the need for extensive tuning.
-Table 4: Effect of Hyperparameter Weights $(\alpha,\beta)$ on MiniGrid Generalization.
-Mean $\pm$ SE of zero-shot solved rates across 12 held-out MiniGrid environments, evaluated under various $(\alpha,\beta)$ settings.
-Except for the final column, TRACED with $(\alpha=1.0,\beta=1.0)$, averaged over five seeds, all other entries report single-seed measurements.
-Bold indicates the best result per task; underline indicates the second-best.
+TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one.
+The ATPL only variant (TRACED - CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED - ATPL) only reaches the second place in two tasks (Hardcore, Roughness).
+Removing both components (the ACCEL baseline) ranks second on one (Stairs).
+These results confirm that both transition-prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
+A.3 Hyperparameter Sensitivity Analysis
+We analyze impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED across the 12 held-out MiniGrid environments and six held-out BipedalWalker terrains.
+Results are reported for a single random seed due to computational constraints.
+Table 5: Effect of hyperparameter weights $(\alpha,\beta)$ on MiniGrid environments.
+Evaluations are performed under varying $(\alpha,\beta)$ settings.
+Bold indicates the best; underline indicates the second-best.
 Environment (0.75, 1.0) (1.25, 1.0) (1.0, 0.75) (1.0, 1.25) (1.25, 0.75) TRACED (1.0, 1.0)
@@
 Mean 0.63 0.45 0.52 0.61 0.67 $\textbf{0.70}\pm 0.04$
-Table 5 reports zero-shot test returns on six held-out BipedalWalker terrains after 10k PPO updates, under five different $(\alpha,\beta)$ settings.
-Notably, three of those configurations outperform all 10k-update baselines (Appendix E).
+In MiniGrid environments, four of the six $(\alpha,\beta)$ settings outperform the 10k-update baselines (Appendix P), and even the weakest setting matches the performance of PLR at 10k (Table 5).
+These results indicate that TRACED provides consistently strong zero-shot generalization across a broad range of hyperparameters, without extensive tuning.
+Table 6: Effect of hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker environments.
+Evaluations are performed under varying $(\alpha,\beta)$ settings.
+Bold indicates the best; underline indicates the second-best.
+Environment (1.25, 0.6) (1.75, 0.6) (1.5, 0.45) (1.5, 0.75) TRACED (1.5, 0.6)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+In BipedalWalker, three of the configurations outperform all 10k-update baselines (Appendix P) (Table 6).
 Even the mean return across all five $(\alpha,\beta)$ choices $(\approx 64.9)$ exceeds every baseline.
 These results demonstrate that TRACED delivers consistently strong zero-shot generalization across a wide range of weight settings, without requiring extensive hyperparameter tuning.
-Table 5: Effect of Hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker Generalization.
-Mean $\pm$ SE of zero-shot test returns on six held-out BipedalWalker terrains, for various $(\alpha,\beta)$ settings.
-Only the final column (TRACED with $(\alpha=1.5,\beta=0.6)$) is averaged over five seeds; all other entries are single-seed runs.
-Bold indicates the best result per terrain; underline indicates the second-best.
-Environment (1.25, 0.6) (1.75, 0.6) (1.5, 0.45) (1.5, 0.75) TRACED (1.5, 0.6)
-Basic 282.06 291.04 281.32 279.37 293.67$\pm$3.56
-Hardcore –7.60 115.87 29.55 50.10 86.83$\pm$17.96
-Stairs –52.78 –21.94 –42.37 –66.02 –29.0$\pm$10.4
-PitGap 12.96 –52.72 –52.21 –122.54 –39.26$\pm$11.42
-Stump –100.95 –40.11 –74.83 –73.93 34.16$\pm$54.58
-Roughness 148.22 187.43 244.57 185.83 193.29$\pm$21.6
-Mean 46.99 79.93 64.34 42.13 89.95$\pm$12.95
-Appendix C Efficiency Analysis
-C.1 Wall-Clock Training Time in BipedalWalker
-Table 6: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration (hours $\pm$ SE over five runs) for each algorithm at 10k and 20k PPO updates.
-Domain TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
-BipedalWalker 35.64 $\pm$ 0.53 33.26 $\pm$ 1.17 70.22 $\pm$ 5.9 21.13 $\pm$ 0.11 41.58 $\pm$ 0.24 35.49 $\pm$ 1.07 71.01 $\pm$ 2.17 18.38 $\pm$ 0.1 37.03 $\pm$ 0.3
-Table 6 reports wall-clock training times on the BipedalWalker domain.
-At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 6% more than ACCEL’s $33.26\pm 1.17$h, yet delivers a 22% increase in mean solved rate (see Section 4).
-C.2 Sample Complexity: Environment Steps
-Table 7: Environment Steps. Total environment interactions (millions) for each method per given number of student PPO updates.
-Environment PPO Updates PLR ADD ACCEL TRACED
-MiniGrid 10k 82M 41M 93M 93M
-MiniGrid 20k 165M 82M 185M
-BipedalWalker 10k 165M 80M 174M 174M
-BipedalWalker 20k 329M 160M 347M
-Table 7 shows total environment steps required to achieve a fixed number of PPO updates.
-TRACED matches ACCEL’s sample complexity exactly, confirming that its superior generalization stems from improved curriculum design rather than additional data.
-Appendix D Visualizing Curriculum Dynamics
+A.4 Ablation Study on the Number of Workers
+We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel.
+The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively.
+To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers.
+Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
+Table 7: Ablation study on Number of Workers. Bold indicates the best; underline indicates the second-best.
+Environment ACCEL 5k ACCEL 10k TRACED 5k TRACED 10k
+16Rooms $0.37\pm 0.17$ $0.78\pm 0.14$ $\textbf{0.95}\pm 0.02$ $\underline{0.93}\pm 0.07$
+16Rooms2 $0.20\pm 0.20$ $\underline{0.77}\pm 0.23$ $0.29\pm 0.14$ $\textbf{0.91}\pm 0.04$
+SimpleCrossing $0.65\pm 0.09$ $0.73\pm 0.07$ $\underline{0.78}\pm 0.02$ $\textbf{0.88}\pm 0.05$
+FourRooms $0.33\pm 0.08$ $\underline{0.45}\pm 0.04$ $0.33\pm 0.04$ $\textbf{0.59}\pm 0.01$
+SmallCorridor $\underline{0.52}\pm 0.21$ $0.27\pm 0.17$ $\textbf{0.62}\pm 0.31$ $0.10\pm 0.04$
+LargeCorridor $\underline{0.42}\pm 0.17$ $0.36\pm 0.30$ $\textbf{0.49}\pm 0.24$ $0.16\pm 0.13$
+Labyrinth $0.15\pm 0.15$ $0.49\pm 0.29$ $\underline{0.67}\pm 0.33$ $\textbf{0.91}\pm 0.09$
+Labyrinth2 $0.0\pm 0.0$ $0.51\pm 0.29$ $\underline{0.65}\pm 0.20$ $\textbf{1.0}\pm 0.0$
+Maze $0.33\pm 0.33$ $0.36\pm 0.32$ $\underline{0.95}\pm 0.04$ $\textbf{0.99}\pm 0.01$
+Maze2 $0.01\pm 0.01$ $\textbf{0.74}\pm 0.23$ $0.22\pm 0.19$ $\underline{0.60}\pm 0.20$
+Maze3 $0.14\pm 0.13$ $0.48\pm 0.27$ $\underline{0.63}\pm 0.29$ $\textbf{1.0}\pm 0.0$
+PerfectMaze(M) $0.20\pm 0.09$ $0.48\pm 0.07$ $\underline{0.58}\pm 0.10$ $\textbf{0.63}\pm 0.13$
+Mean $0.28\pm 0.10$ $0.54\pm 0.09$ $\underline{0.60}\pm 0.06$ $\textbf{0.72}\pm 0.09$
+Appendix B Long-term Analysis on TRACED
+Maze
+FourRooms
+Labyrinth
+PerfectMaze(M)
+SimpleCrossing
+SmallCorridor
+Figure 12: TRACED Solved-Rate Time Series. Solved rate progression on MiniGrid tasks plotted over 0-45k PPO updates.
+Maze
+FourRooms
+Labyrinth
+PerfectMaze(M)
+SimpleCrossing
+SmallCorridor
+Figure 13: ACCEL Solved-Rate Time Series. Solved rate progression on MiniGrid tasks plotted over 0-45k PPO updates.
+To study long-horizon behavior, we ran single-seed, 45k PPO updates for both TRACED and ACCEL in the MiniGrid environment.
+Both methods exhibit post-convergence oscillations in success rate (i.e., solved rate).
+TRACED reaches near-peak success rapidly, but its performance subsequently oscillates rather than remaining perfectly stable (Figure 12).
+The same post-peak fluctuations appear under the ACCEL curriculum (Figure 13): although ACCEL attains its maximum more slowly, it shows a similar up-and-down pattern thereafter.
+These oscillations likely reflect inherent instability of the RL policy on held-out levels, and may be further influenced by ACCEL’s mutation dynamics, rather than a pathology introduced by our task-prioritization strategy.
+For this reason, we report TRACED’s performance at 10k updates (rather than 20k), before long-horizon oscillations confound comparisons.
+Mitigating these fluctuations in the student policy is an important direction for future work.
+Appendix C Visualizing Curriculum Dynamics
 (a) 2.5k, Low ATPL, Low Co-Learnability
@@
 Figure 14: Representative terrains from BipedalWalker selected by ATPL and co-learnability at two training stages.
-The top two rows show terrains after 2.5k PPO updates (first row: low ATPL; second row: high ATPL), and the bottom two rows after 5k updates (third row: low ATPL; fourth row: high ATPL).
-Within each row, the left column displays low co-learnability examples and the right column high co-learnability examples.
-Figure 12 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at two training milestones (2.5k and 5k PPO updates).
+The top two rows show terrains after 2.5k PPO updates, and the bottom two rows after 5k updates.
+Figure 14 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at 2.5k and 5k PPO updates.
 ATPL alone captures the raw challenge level: low-ATPL terrains (first and third rows) remain relatively smooth, while high-ATPL terrains (second and fourth rows) feature larger gaps and steeper bumps.
@@
 Together, these visualizations confirm that ATPL drives a steadily increasing complexity trajectory, while Co-Learnability smooths it to sustain robust, transferable learning throughout the curriculum.
-Appendix E Numerical Results
-E.1 Minigrid Environment
-Table 8: MiniGrid Zero-Shot Solved Rates.
-Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks for DR, PLR, ADD, ACCEL, and TRACED, evaluated after 10k and 20k PPO updates over 10 seeds.
-Bold indicates the best result per task; underline indicates the second-best.
-Environment Update DR PLR ADD ACCEL TRACED
-16Rooms 10k 0.4 $\pm$ 0.12 0.81 $\pm$ 0.06 0.72 $\pm$ 0.1 0.76 $\pm$ 0.12 0.53 $\pm$ 0.14
-20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
-16Rooms2 10k 0.1 $\pm$ 0.09 0.42 $\pm$ 0.11 0.14 $\pm$ 0.05 0.48 $\pm$ 0.13 0.58 $\pm$ 0.13
-20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
-SimpleCrossing 10k 0.58 $\pm$ 0.05 0.72 $\pm$ 0.03 0.63 $\pm$ 0.03 0.81 $\pm$ 0.03 0.85 $\pm$ 0.03
-20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
-FourRooms 10k 0.32 $\pm$ 0.04 0.47 $\pm$ 0.01 0.39 $\pm$ 0.03 0.46 $\pm$ 0.02 0.46 $\pm$ 0.02
-20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
-SmallCorridor 10k 0.3 $\pm$ 0.07 0.52 $\pm$ 0.13 0.47 $\pm$ 0.1 0.37 $\pm$ 0.14 0.49 $\pm$ 0.13
-20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
-LargeCorridor 10k 0.36 $\pm$ 0.13 0.33 $\pm$ 0.13 0.35 $\pm$ 0.09 0.44 $\pm$ 0.15 0.43 $\pm$ 0.11
-20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
-Labyrinth 10k 0.1 $\pm$ 0.1 0.5 $\pm$ 0.15 0.22 $\pm$ 0.13 0.6 $\pm$ 0.15 0.93 $\pm$ 0.03
-20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
-Labyrinth2 10k 0.1 $\pm$ 0.1 0.28 $\pm$ 0.13 0.0 $\pm$ 0.0 0.52 $\pm$ 0.14 0.96 $\pm$ 0.03
-20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
-Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
-20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
-Maze2 10k 0.19 $\pm$ 0.12 0.43 $\pm$ 0.14 0.09 $\pm$ 0.08 0.44 $\pm$ 0.15 0.6 $\pm$ 0.1
-20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
-Maze3 10k 0.27 $\pm$ 0.13 0.57 $\pm$ 0.13 0.2 $\pm$ 0.1 0.78 $\pm$ 0.13 0.81 $\pm$ 0.06
-20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
-PerfectMaze(M) 10k 0.06 $\pm$ 0.02 0.24 $\pm$ 0.05 0.17 $\pm$ 0.08 0.41 $\pm$ 0.05 0.64 $\pm$ 0.07
-20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
-Mean 10k 0.23 $\pm$ 0.05 0.45 $\pm$ 0.04 0.28 $\pm$ 0.04 0.54 $\pm$ 0.03 0.66 $\pm$ 0.03
-20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
-Table 9: MiniGrid Ablation Results.
-Mean $\pm$ SE of zero-shot solved rates on 12 held-out MiniGrid tasks, comparing ACCEL (10k), TRACED – ATPL (10k), TRACED – CL (10k), and TRACED (10k) over 5 seeds.
-For TRACED and ACCEL, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation.
-Bold indicates the best result per task; underline indicates the second-best.
-Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
-16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
-16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
-SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
-FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
-SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
-LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
-Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
-Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
-Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
-Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
-Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
-PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
-Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
-Table 10: PerfectMaze Zero-Shot Solved Rates.
-Mean $\pm$ SE of zero-shot success rates on PerfectMazeLarge and PerfectMazeXL after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, and TRACED over 10 seeds.
-Bold indicates the best result per environment; underline indicates the second-best.
-Environment Update DR PLR ADD ACCEL TRACED
-PerfectMaze(Large) 10k 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01 0.04 $\pm$ 0.04 0.09 $\pm$ 0.03 0.27 $\pm$ 0.07
-20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
-PerfectMaze(XL) 10k 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.02 0.01 $\pm$ 0.01 0.1 $\pm$ 0.04
-20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
-Table 8 reports per-task zero-shot solved rates on the 12 held-out MiniGrid mazes, following [21].
-We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (at both 10k and 20k).
-TRACED at 10k achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best rate in one maze (Maze2).
-Averaged across all 12 mazes, TRACED attains a mean solved rate of 0.66, outperforming every other baseline configuration.
-Table 9 presents zero-shot solved rates on 12 held-out MiniGrid mazes for TRACED and its ablations at 10k updates.
-Full TRACED attains the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the other six, with a mean solved rate of 0.70, outperforming every variant.
-These results confirm that both transition prediction error and Co-Learnability are critical to TRACED’s superior, consistent performance across diverse tasks.
-Table 10 compares zero-shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large-scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL.
-After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third-highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k.
-These results underscore TRACED’s ability to scale to extreme maze sizes.
-E.2 BipedalWalker Environment
-Table 11 compares TRACED’s zero-shot returns against DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains.
-At 10k updates, TRACED attains the highest mean return in three environments (Basic, Hardcore, and Stump) and achieves an overall mean of 89.95 $\pm$ 31.72, outperforming all baselines.
-Table 11: BipedalWalker Zero-Shot Test Returns.
-Mean $\pm$ SE of episode returns on six held-out BipedalWalkerHardcore terrains after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, ACCEL-CENIE, and TRACED over 5 seeds.
-Bold indicates the best result per terrain; underline indicates the second-best.
-ACCEL-CENIE is estimated by applying the original CENIE paper’s ACCEL-to-ACCEL-CENIE performance ratio [20] to our ACCEL implementation.
-Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
-Basic 10k 112.56 $\pm$ 62.99 131.63 $\pm$ 56.58 119.43 $\pm$ 63.72 281.65 $\pm$ 5.25 273.56 $\pm$ 1.5 293.67 $\pm$ 3.56
-20k 68.17 $\pm$ 44.71 97.67 $\pm$ 61.73 63.64 $\pm$ 62.71 281.85 $\pm$ 3.72 275.04 $\pm$ 3.19
-Hardcore 10k -19.67 $\pm$ 7.35 -17.36 $\pm$ 5.99 3.83 $\pm$ 9.07 37.59 $\pm$ 15.0 66.83 $\pm$ 17.48 86.83 $\pm$ 17.96
-20k -16.98 $\pm$ 5.8 -23.7 $\pm$ 2.05 10.48 $\pm$ 20.12 59.23 $\pm$ 25.5 84.09 $\pm$ 34.32
-Stairs 10k -17.6 $\pm$ 7.21 -7.23 $\pm$ 7.16 -5.87 $\pm$ 9.07 -38.71 $\pm$ 10.54 -30.4 $\pm$ 8.35 -29.0 $\pm$ 10.4
-20k -9.11 $\pm$ 7.67 -10.2 $\pm$ 4.5 -0.66 $\pm$ 9.2 -46.34 $\pm$ 5.78 -36.39 $\pm$ 11.48
-PitGap 10k -34.39 $\pm$ 15.66 -46.21 $\pm$ 15.69 -17.8 $\pm$ 13.82 -65.07 $\pm$ 7.57 -39.81 $\pm$ 18.75 -39.26 $\pm$ 11.42
-20k -26.03 $\pm$ 10.98 -48.72 $\pm$ 12.23 -7.65 $\pm$ 7.52 -64.89 $\pm$ 18.93 -45.92 $\pm$ 33.9
-Stump 10k -24.99 $\pm$ 6.76 -27.74 $\pm$ 6.33 -26.15 $\pm$ 8.49 -79.18 $\pm$ 5.45 -60.05 $\pm$ 10.96 34.16 $\pm$ 54.58
-20k -25.52 $\pm$ 12.39 -23.61 $\pm$ 2.86 3.26 $\pm$ 19.1 -67.18 $\pm$ 15.56 -46.34 $\pm$ 176.21
-Roughness 10k -0.66 $\pm$ 12.01 3.55 $\pm$ 9.91 20.61 $\pm$ 13.78 161.72 $\pm$ 28.36 174.92 $\pm$ 2.25 193.29 $\pm$ 21.6
-20k 0.19 $\pm$ 8.18 -2.73 $\pm$ 9.51 18.71 $\pm$ 21.59 213.48 $\pm$ 7.69 224.4 $\pm$ 8.63
-Mean 10k 2.54 $\pm$ 15.44 6.1 $\pm$ 10.69 14.4 $\pm$ 14.05 49.67 $\pm$ 11.24 64.18 89.95 $\pm$ 12.95
-20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
-Table 12 compares zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED.
-A trajectory is considered “solved” if its return exceeds 230.
-At 10k updates, TRACED achieves the top solved rate in five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36—substantially higher than ACCEL’s 0.29 and the other baselines.
-No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
-Table 12: BipedalWalker Zero-Shot Solved Rates.
-Mean $\pm$ SE of solved rates, defined as the fraction of episodes exceeding a return of 230, on 6 held-out BipedalWalkerHardcore terrains (Basic, Hardcore, Stairs, PitGap, Stump, Roughness) after 10k and 20k PPO updates.
-Results compare DR, PLR, ADD, ACCEL, and TRACED over 5 seeds.
-Bold indicates the best per terrain; underline indicates the second-best.
-Environment Update DR PLR ADD ACCEL TRACED
-Basic 10k 0.35 $\pm$ 0.2 0.43 $\pm$ 0.21 0.34 $\pm$ 0.21 0.98 $\pm$ 0.02 1.00 $\pm$ 0.0
-20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
-Hardcore 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.01 $\pm$ 0.01 0.15 $\pm$ 0.03 0.28 $\pm$ 0.05
-20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
-Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
-20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
-PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
-20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
-Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
-20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
-Roughness 10k 0.01 $\pm$ 0.01 0.0 $\pm$ 0.01 0.01 $\pm$ 0.01 0.56 $\pm$ 0.1 0.65 $\pm$ 0.08
-20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
-Mean 10k 0.06 $\pm$ 0.04 0.07 $\pm$ 0.04 0.06 $\pm$ 0.04 0.29 $\pm$ 0.02 0.36 $\pm$ 0.03
-20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03
-Appendix F Implementation Details
-F.1 Minigrid Environment
+Appendix D Level Evolution
+D.1 Visualization of Level Evolution in Minigrid
+The visualization of how the Minigrid environment evolves as the number of blocks increases (Figure 15).
+Each step of the evolutionary process produces an edited level that has a high learning efficiency.
+Figure 15: Visualization of a single level’s evolving progression in the MiniGrid environment.
+Starting from top-left, ending bottom-right.
+This progress is automatically designed by our TRACED algorithm.
+D.2 Evolution of Levels in BipedalWalker
+Figure 16 shows the complexity metric results trained by three methods, TRACED (Ours), ACCEL + ATPL, ACCEL.
+Starting with plain terrain (near zero point), all three methods guide levels with more complex terrain.
+With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL.
+This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
+(a) Stump height low
+(b) Stump height high
+(c) Stair height low
+(d) Stair height high
+(e) Stair height step
+(f) Pit gap low
+(g) Pit gap high
+(h) Roughness
+Figure 16: Emergent BipedalWalker terrain complexity metrics.
+Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue).
+The result is averaged over five random seeds.
+This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
+D.3 Visualization of Level Evolution in BipedalWalker
+Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases.
+The process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
+Figure 17: Visualization of the level evolving progression in the BipedalWalker environment.
+Starting from top-left, ending bottom-right.
+in this example, starting with plain terrain, the pits are created and their number increases, then the roughness increases, then stairs and stumps are created and their number, width, and height increase.
+This progress is automatically designed by our TRACED algorithm.
+Appendix E Agent Trajectory Visualizations Across Environments
+E.1 MiniGrid
+Figure 18: Agent trajectory visualization in the PerfectMaze evaluation task.
+TRACED enables efficient path planning and robust generalization to challenging maze environments, as demonstrated by the agent’s successful navigation to the goal (green).
+E.2 BipedalWalker
+Figure 19: Agent trajectory visualization in the BipedalWalkerHardcore.
+TRACED agent successfully overcomes various combined obstacles, including stairs, pit gaps, stumps, and roughness, demonstrating robust generalization to complex environments.
+Appendix F Efficiency Analysis
+F.1 Wall-Clock Training Time in BipedalWalker
+Table 8 reports wall-clock training times on the BipedalWalker domain, in hours (mean $\pm$ s.e.) over five runs at 10k and 20k PPO updates.
+At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 7% more than ACCEL’s $33.26\pm 1.17$ h, yet achieves a 22% higher mean solved rate (Section 4).
+Table 8: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration for each algorithm
+Domain TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
+BipedalWalker 35.64 $\pm$ 0.53 33.26 $\pm$ 1.17 70.22 $\pm$ 5.9 21.13 $\pm$ 0.11 41.58 $\pm$ 0.24 35.49 $\pm$ 1.07 71.01 $\pm$ 2.17 18.38 $\pm$ 0.1 37.03 $\pm$ 0.3
+F.2 Sample Complexity: Environment Steps
+Table 9 shows total environment steps required to achieve a fixed number of PPO updates.
+TRACED matches ACCEL’s sample complexity exactly, confirming that its superior generalization stems from improved curriculum design rather than additional data.
+Table 9: Environment Steps. Total environment interactions (millions) for each method per given number of student PPO updates.
+Environment PPO Updates PLR ADD ACCEL TRACED
+MiniGrid 10k 82M 41M 93M 93M
+MiniGrid 20k 165M 82M 185M
+BipedalWalker 10k 165M 80M 174M 174M
+BipedalWalker 20k 329M 160M 347M
+Appendix G Overall Workflow
+Algorithm 1 summarizes our overall UED procedure under the TRACED framework.
+We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority.
+At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
+$\widehat{\mathrm{Regret}}(\tau)\;=\;\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau).$
+If the buffer is full, we evict the lowest‐priority task before appending.
+Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret.
+We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer.
+Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq.
+7) before proceeding to the next timestep.
+Algorithm 1 UED Workflow with Transition‐Prediction Loss and Co‑Learnability
+1:Given: policy $\pi_{\phi}$, level generator $\mathcal{G}$, buffer capacity $K$, replay probability $p_{\mathrm{replay}}$, batch size $B$, mutation count $N_{\text{mutate}}$, scaling factors $\alpha,\beta$
+2:Initialize Task Buffer $\Lambda\leftarrow\emptyset$, timestep $t\leftarrow 0$
+3:while policy has not converged do
+4:  Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
+[…]
+6:   Sample new task $\theta\sim\mathcal{G}$
+7:   Collect trajectory $\tau$ by executing $\pi_{\phi}$ on $\theta$ (no gradient)
+8:   Compute regret estimate
+$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)+\alpha\,\mathrm{ATPL}(\tau)$
+9:   Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest‐priority)
+[…]
+11:   Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\frac{1}{\mathrm{TaskPriority}(\cdot,t)}$
+12:   for $k=1\ldots B$ do
+13:     Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
+14:     Update policy $\phi$ using rewards from $\tau_{k}$
+15:     Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
+16:   end for
+17:   Select the $N$ tasks $\{\theta_{k^{\prime}}\}$ with smallest $\widehat{\mathrm{Regret}}(\tau_{k^{\prime}})$
+18:   for each selected $\theta_{k^{\prime}}$ do
+19:     $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
+20:     Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
+[…]
+22:     Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
+23:   end for
+24:  end if
+25:  Buffer update:
+26:  for each task $i\in\Lambda$ updated at $t$ do
+27:   Recompute $\mathrm{TaskDifficulty}(i,t)$
+28:   Recompute $\mathrm{CoLearnability}(i,t-1)$
+29:   Update $\mathrm{TaskPriority}(i,t)$ via Eq. 7
+30:  end for
+31:  $t\leftarrow t+1$
+32:end while
+Appendix H Implementation Details
+H.1 Minigrid Environment
 SixteenRooms
@@
 In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths.
-SimpleCrossing and FourRooms are adapted from MiniGrid [16], and PerfectMaze from REPAIRED [6].
-The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid [16].
+SimpleCrossing and FourRooms are adapted from MiniGrid (Chevalier-Boisvert et al., 2023), and PerfectMaze from REPAIRED (Jiang et al., 2021a).
+The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (Chevalier-Boisvert et al., 2023).
 Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself.
@@
 At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction.
-The action space is defined as a discrete set of seven possible actions, although only three—turn left, turn right, and move forward—are relevant to maze navigation.
+The action space is defined as a discrete set of seven possible actions, although only three, turn left, turn right, and move forward, are relevant to maze navigation.
 These actions are selected without masking the unused action outputs.
@@
 If the agent fails to reach the goal within the time limit, it does not receive a reward.
-To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work [4, 6].
+To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (Dennis et al., 2020; Jiang et al., 2021a).
 Figure 20 displays all evaluation tasks.
-F.2 BipedalWalker Environment
-The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym [23].
+H.2 BipedalWalker Environment
+The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman et al., 2016).
 At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead.
@@
 These obstacles are procedurally generated with randomized parameters.
-The stair height is randomly set to either $+1$ (ascending) or $-1$ (descending), with the number of steps and step width randomly sampled between 3–5 steps and 4–5 units, respectively.
-The pit width is determined by sampling a single integer between 3–5 units.
-Both the height and width of each stump are sampled as a single random integer between 1–3 units, ensuring the stumps are always square-shaped.
+The stair height is randomly set to either +1+1 (ascending) or 1-1 (descending), with the number of steps and step width randomly sampled between 3-5 steps and 4-5 units, respectively.
+The pit width is determined by sampling a single integer between 3-5 units.
+Both the height and width of each stump are sampled as a single random integer between 1-3 units, ensuring the stumps are always square-shaped.
 Terrain roughness noise range is the same as BipedalWalker-v3.
@@
 Figure 21: Evaluation Terrains in BipedalWalker.
-Six held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
-F.3 Network Structures
+6 held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
+H.3 Network Structures
 For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1.
-an image encoder that compresses the input observation $s_{t}$ into a 128-dimensional latent embedding, 2.
+an image encoder that compresses the input observation $s_{t}$ into a 128‑dimensional latent embedding, 2.
 an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3.
 and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output.
-We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video-prediction tasks under partial observability [25, 26].
+We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (Ha & Schmidhuber, 2018; Oh et al., 2015).
 For walking in challenging terrain (BipedalWalker), the model instead concatenates the raw state vector $s_{t}$ with $a_{t}$ and feeds this into an LSTM (hidden size 128, 1 layer); the recurrent output is then linearly mapped to the predicted next state $\hat{s}_{t+1}$.
@@
 $\ell_{t}=\begin{cases}\|\,\hat{s}_{t+1}-s_{t+1}\|_{1},&\text{in MiniGrid environments},\\ \|\hat{s}_{t+1}-s_{t+1}\|_{2}^{2},&\text{in BipedalWalker}.\end{cases}$
-We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction [27], while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains [28].
+We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (Zhao et al., 2016), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (Mnih et al., 2015).
+We adopt ACCEL’s network architecture (Parker-Holder et al., 2022), differing only in the LSTM design choice.
 The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment.
@@
 The PPO learning process optimizes a clipped surrogate policy loss with clip parameter $\epsilon$ = 0.2 combined with a value function loss.
-Appendix G Overall Workflow
-Algorithm 1 summarizes our overall UED procedure under the TRACED framework.
-At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
-$\widehat{\mathrm{Regret}}(\tau)\;=\;\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau).$
-If the buffer is full, we evict the lowest-priority task before appending.
-Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret.
-We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer.
-Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq.
-(7)) before proceeding to the next timestep.
-Algorithm 1 UED Workflow with Transition-Prediction Loss and Co-Learnability
-1:Given: policy $\pi_{\phi}$, level generator $\mathcal{G}$, buffer capacity $K$, replay probability $p_{\mathrm{replay}}$, batch size $B$, mutation count $N_{\text{mutate}}$, scaling factors $\alpha,\beta$
-2:Initialize Task Buffer $\Lambda\leftarrow\emptyset$, timestep $t\leftarrow 0$
-3:while policy has not converged do
-4:     Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
-5:     if $d_{t}=0$ $\triangleright$ Exploration phase
-6:         Sample new task $\theta\sim\mathcal{G}$
-7:         Collect trajectory $\tau$ by executing $\pi_{\phi}$ on $\theta$ (no gradient)
-8:         Compute regret estimate
-$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)+\alpha\,\mathrm{ATPL}(\tau)$
-9:         Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest-priority)
-10:     $\triangleright$ Replay + Mutation phase
-11:         Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\mathrm{TaskPriority}(\cdot,t)$
-12:         for $k=1\ldots B$ do
-13:              Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
-14:              Update policy $\phi$ using rewards from $\tau_{k}$
-15:              Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
-16:         end for
-17:         Select the $N$ tasks $\{\theta_{k^{\prime}}\}$ with smallest $\widehat{\mathrm{Regret}}(\tau_{k^{\prime}})$
-18:         for each selected $\theta_{k^{\prime}}$ do
-19:              $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
-20:              Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
-21:              Compute $\widehat{\mathrm{Regret}}(\tilde{\tau})$
-22:              Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
-23:         end for
-24:     end if
-25:     Buffer update:
-26:     for each task $i\in\Lambda$ updated at $t$ do
-27:         Recompute $\mathrm{TaskDifficulty}(i,t)$
-28:         Recompute $\mathrm{CoLearnability}(i,t-1)$
-29:         Update $\mathrm{TaskPriority}(i,t)$ via Eq. (7)
-30:     end for
-31:     $t\leftarrow t+1$
-32:end while
-Appendix H Hyperparameters
-Table 13 contains the hyperparameters we use.
-Hyperparameter values were adopted directly from the DCD repository [29].
-We perform hyperparameter tuning in MiniGrid and Bipedal environments, focusing on two key components, $\alpha$ and $\beta$.
-We explored values in the ranges $\alpha\in[0.5,1.0,1.5]$, $\beta\in[0.6,0.8,1.0]$.
+Appendix I Difference with Former Studies
+I.1 Difference with ACCEL
+Our approach builds on ACCEL’s evolutionary, regret-based curriculum design, refining both its regret estimator and its sampling criterion.
+Like ACCEL, we maintain a level buffer and record each task’s difficulty.
+However, instead of relying solely on Positive Value Loss (PVL) to approximate regret, we decompose regret into value error and transition-prediction error and incorporate both components into our difficulty measure.
+Furthermore, prior UED methods sample levels independently based solely on their individual difficulty, ignoring cross-task transfer effects.
+We address this with our Co-Learnability metric, which quantifies how training on one task influences regret on others by measuring the average change in their difficulty after replay.
+By integrating Co-Learnability into our priority score, we can prioritize tasks that not only challenge the agent but also deliver broad transfer benefits across the entire task space.
+I.2 Difference with CENIE
+CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL.
+It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state-action space.
+This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
+However, CENIE’s diversity-driven objective differs from ours.
+TRACED’s primary goal is to approximate task difficulty more faithfully, by decomposing regret into value and transition-prediction errors, and to capture how training on one task transfers to others via our Co-Learnability metric.
+While CENIE accounts for per-task novelty and regret, it ignores cross-task transfer effects.
+Moreover, TRACED is agnostic to the choice of level generator or editor, meaning our method could be combined with CENIE’s novelty scoring to yield a curriculum that balances both transfer-aware difficulty and environmental diversity.
+Appendix J Related Works
+Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (Wang et al., 2019; 2020; Dennis et al., 2020; Chung et al., 2024).
+PAIRED (Dennis et al., 2020) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
+However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula.
+To mitigate these limitations, Mediratta et al.
+(2023) introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (Azad et al., 2023a) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
+Another line of work is based on Prioritized Level Replay (PLR) (Jiang et al., 2021b), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
+This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent.
+PLR  (Jiang et al., 2021a) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
+This leads to better performance on unseen environments counterintuitively.
+ACCEL (Parker-Holder et al., 2022) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
+This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks.
+However, these methods (Jiang et al., 2021b; a; Parker-Holder et al., 2022) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
+Moving beyond regret-based approaches, CENIE (Teoh et al., 2024) integrates novelty with regret to expose agents to more diverse learning situations.
+Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer.
+This curriculum-awareness leads to better exploration and generalization.
+DIPLR (Li et al., 2023) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories.
+MBeDED (Li et al., 2024) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
+It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task.
+So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
+UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains.
+For Multi-Agent Reinforcement Learning, MAESTRO (Samvelyan et al., 2023) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
+In contrast, RACCOON (Erlebach & Cook, 2024) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
+For meta-RL, GROOVE (Jackson et al., 2023) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
+DIVA (Costales & Nikolaidis, 2024) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
+In sim-to-real settings, GCL (Wang et al., 2024) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
+Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (Narvekar et al., 2020).
+Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (Klink et al., 2020).
+Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (Eimer et al., 2021; Klink et al., 2022).
+WAKER (Rigter et al., 2024) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
+However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
+Appendix K More Details On Baseline Algorithms
+In Unsupervised Environment Design (UED), Domain Randomization (DR) constructs the curriculum by uniformly sampling tasks from the environment parameter space $\Theta$.
+Formally, DR samples tasks according to $\theta\sim p(\Theta)$ where $p(\Theta)$ represents a uniform distribution over the task parameter space.
+The agent is trained on these randomly sampled tasks.
+Robust PLR (PLR(Jiang et al., 2021a) was introduced as an extension of the original PLR (Jiang et al., 2021b).
+While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay.
+When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks.
+The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer.
+Policy updates are exclusively performed on tasks sampled from the replay buffer.
+By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
+Adversarially Compounding Complexity by Editing Levels (ACCEL) (Parker-Holder et al., 2022) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
+This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret.
+ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation.
+The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities.
+The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid.
+This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity.
+Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
+Coverage-based Evaluation of Novelty In Environment (CENIE) (Teoh et al., 2024) introduces environment novelty as a complementary objective to regret in UED.
+CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum.
+The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space.
+At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences.
+Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution.
+The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state-action pairs and the learned distribution of past state-action experiences.
+CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
+Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung et al., 2024) frames environment generation as a diffusion process steered by the agent’s regret.
+Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
+In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (Jiang et al., 2022), and assessed ADD with the ADD codebase (Chung et al., 2024).
+CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
+Appendix L Limitations
+As with any research, our approach presents several limitations that highlight opportunities for future investigation:
+Co-Learnability.
+In this paper, we proposed Co-Learnability, simple yet effective method.
+Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
+ATPL. In environments with highly stochastic transitions, the ATPL signal (transition-prediction error) can become noisy, potentially reducing the fidelity of our regret approximation.
+Sampling levels.
+We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods.
+However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
+Heuristic weighting.
+We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty.
+While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
+Extension to other RL algorithms.
+Our experiments have been conducted with PPO-based student agents.
+While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum.
+Off-policy methods such as SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
+Moreover, Model-based approaches like DreamerV3 (Hafner et al., 2025) build world models for planning, offering complementary insights into curriculum design.
+Evaluating these methods under our curriculum framework remains an important direction for future work.
+Extend to other UED methods.
+We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost.
+While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (Jiang et al., 2021a) and CENIE (Teoh et al., 2024)) and conduct comprehensive experiments to assess its performance in these settings.
+Evaluation in other domains.
+Our experimental validation has been conducted in MiniGrid (Chevalier-Boisvert et al., 2023; Dennis et al., 2020) and BipedalWalker (Brockman et al., 2016; Parker-Holder et al., 2022).
+Evaluating TRACED in other domains such as Car Racing (Brockman et al., 2016; Jiang et al., 2021a), physical robotics applications (Mahmood et al., 2018), and reasoning task ARC (Chollet, 2019) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
+Appendix M Experimental Setup and Reproducibility
+All experiments were conducted using Python 3.8.20 on Ubuntu 22.04.4 LTS.
+The hardware setup included an AMD EPYC 7543 32-core processor, 80GB of RAM, and an NVIDIA A100 GPU.
+Experiments were run on a shared server with 8 GPUs, with each experiment using a single GPU.
+We implemented all experiments in Python 3.8.20, using PyTorch(v1.9.0+cu111).
+The environment stack was built using MiniGrid (v1.0.1), Gym (v0.15.7).
+Appendix N LLM Usage
+We used a large language model (LLM) solely for language editing.
+Concretely, the LLM assisted with grammar and style polishing, LaTeX phrasing (e.g., equation and caption wording), and improving clarity and concision of author-written text.
+The LLM was not used to generate ideas, design algorithms, select hyperparameters, run experiments, analyze data, create figures/tables, write code, or produce mathematical results.
+Appendix O Hyperparameters
+Table 10 summarizes the hyperparameters used.
+Unless otherwise noted, values are adopted directly from the DCD repository (Jiang et al., 2022).
+The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4).
+We tune two key weights, $\alpha$ and $\beta$, on MiniGrid and BipedalWalker, exploring $\alpha\in\{0.5,1.0,1.5\}$ and $\beta\in\{0.6,0.8,1.0\}$.
 Table 10: Hyperparameters used for training each method in each environment.
@@
 Staleness coefficient, $\rho$ 0.3 0.5
-Appendix I Level Evolution
-I.1 Visualization of Level Evolution in Minigrid
-The visualization of how the Minigrid environment evolves as the number of blocks increases is shown in Figure 15.
-Each step of the evolutionary process produces an edited level that has a high learning efficiency.
-Figure 15: Visualization of a single level’s evolving progression in the MiniGrid environment.
-Starting from top-left, ending bottom-right.
-This progress is automatically designed by our TRACED algorithm.
-I.2 Evolution of Levels in BipedalWalker
-Figure 16 shows the complexity metric results trained by 3 methods, TRACED (Ours), ACCEL + ATPL, ACCEL.
-Starting with plain terrain (near zero point), all 3 methods guide levels with more complex terrain.
-With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL.
-This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
-(a) Stump height low
-(b) Stump height high
-(c) Stair height low
-(d) Stair height high
-(e) Stair height step
-(f) Pit gap low
-(g) Pit gap high
-(h) Roughness
-Figure 16: Emergent BipedalWalker terrain complexity metrics.
-Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue).
-The result is averaged over 5 random seeds.
-This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
-I.3 Visualization of Level Evolution in BipedalWalker
-Figure 17: Visualization of the level evolving progression in the BipedalWalker environment.
-Starting from top-left, ending bottom-right.
-in this example, starting with plain terrain, the pits are created and their number increases, then the roughness increases, then stairs and stumps are created and their number, width, and height increase.
-This progress is automatically designed by our TRACED algorithm.
-Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases.
-As we can see, the process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
-Appendix J Agent Trajectory Visualizations Across Environments
-J.1 MiniGrid
-Figure 18 shows a representative rollout in the PerfectMaze environment.
-Figure 18: Agent trajectory visualization in the PerfectMaze evaluation task.
-TRACED enables efficient path planning and robust generalization to challenging maze environments, as demonstrated by the agent’s successful navigation to the goal (green).
-J.2 BipedalWalker
-Figure 19 illustrates a zero-shot rollout in BipedalWalkerHardcore.
-The agent dynamically modulates joint torques to ascend and descend steps, leap over gaps, step onto stumps, and maintain balance on uneven ground, confirming TRACED’s effectiveness in teaching transferrable locomotion skills.
-Figure 19: Agent trajectory visualization in the BipedalWalkerHardcore.
-TRACED agent successfully overcomes various combined obstacles, including stairs, pit gaps, stumps, and roughness, demonstrating robust generalization to complex environments.
-Appendix K Difference with Former Studies
-K.1 Difference with ACCEL
-Our approach builds on ACCEL’s evolutionary, regret-based curriculum design, refining both its regret estimator and its sampling criterion.
-Like ACCEL, we maintain a level buffer and record each task’s difficulty.
-However, instead of relying solely on Positive Value Loss (PVL) to approximate regret, we decompose regret into value error and transition-prediction error and incorporate both components into our difficulty measure.
-Furthermore, prior UED methods sample levels independently based solely on their individual difficulty, ignoring cross-task transfer effects.
-We address this with our Co-Learnability metric, which quantifies how training on one task influences regret on others by measuring the average change in their difficulty after replay.
-By integrating Co-Learnability into our priority score, we can prioritize tasks that not only challenge the agent but also deliver broad transfer benefits across the entire task space.
-K.2 Difference with CENIE
-CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL.
-It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state–action space.
-This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
-However, CENIE’s diversity-driven objective differs from ours.
-TRACED’s primary goal is to approximate task difficulty more faithfully, by decomposing regret into value and transition-prediction errors, and to capture how training on one task transfers to others via our Co-Learnability metric.
-While CENIE accounts for per-task novelty and regret, it ignores cross-task transfer effects.
-Moreover, TRACED is agnostic to the choice of level generator or editor, meaning our method could be combined with CENIE’s novelty scoring to yield a curriculum that balances both transfer-aware difficulty and environmental diversity.
-Appendix L Related Work
-Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments [24, 30, 4, 19].
-PAIRED [4] trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
-However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula.
-To mitigate these limitations, Mediratta et al.
-[9] introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr [31] employs variational autoencoder–based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
-Another line of work is based on Prioritized Level Replay (PLR) [5], which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
-This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent.
-PLR  [6] make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
-This leads to better performance on unseen environments counterintuitively.
-ACCEL [7] extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
-This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks.
-However, these methods [5, 6, 7] approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
-Moving beyond regret-based approaches, CENIE [20] integrates novelty with regret to expose agents to more diverse learning situations.
-Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer.
-This curriculum-awareness leads to better exploration and generalization.
-DIPLR [32] introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state–action trajectories.
-MBeDED [33] introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
-It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task.
-So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
-UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains.
-For Multi-Agent Reinforcement Learning, MAESTRO [34] extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
-In contrast, RACCOON [10] applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
-For meta-RL, GROOVE [35] suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
-DIVA [36] is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
-In sim-to-real settings, GCL [37] aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
-Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task [38].
-Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress [39].
-Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process [40, 41].
-WAKER [42] leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
-However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
-Appendix M More Details On Baseline Algorithms
-In Unsupervised Environment Design (UED), Domain Randomization (DR) constructs the curriculum by uniformly sampling tasks from the environment parameter space $\Theta$.
-Formally, DR samples tasks according to $\theta\sim p(\Theta)$ where $p(\Theta)$ represents a uniform distribution over the task parameter space.
-The agent is trained on these randomly sampled tasks.
-Robust PLR (PLR[6] was introduced as an extension of the original PLR [5].
-While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay.
-When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks.
-The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer.
-Policy updates are exclusively performed on tasks sampled from the replay buffer.
-By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
-Adversarially Compounding Complexity by Editing Levels (ACCEL) [7] actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
-This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret.
-ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation.
-The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities.
-The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid.
-This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity.
-Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
-Coverage-based Evaluation of Novelty In Environment (CENIE) [20] introduces environment novelty as a complementary objective to regret in UED.
-CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum.
-The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space.
-At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences.
-Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution.
-The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state–action pairs and the learned distribution of past state–action experiences.
-CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
-Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) [19] frames environment generation as a diffusion process steered by the agent’s regret.
-Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
-In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase [29], and assessed ADD with the ADD codebase [19].
-CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
-Appendix N Long-term Analysis on TRACED
-Maze
-FourRooms
-Labyrinth
-PerfectMaze(M)
-SimpleCrossing
-SmallCorridor
-Figure 20: TRACED Solved-Rate Time Series.
-Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed.
-Each curve shows the instantaneous solved rate at each update step.
-Maze
-FourRooms
-Labyrinth
-PerfectMaze(M)
-SimpleCrossing
-SmallCorridor
-Figure 21: ACCEL Solved-Rate Time Series.
-Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed.
-Each curve shows the instantaneous solved rate at each update step.
-TRACED reaches near-peak solved rates very quickly.
-However, its performance subsequently oscillates rather than remaining perfectly stable (Figure 20).
-Crucially, this same post-peak fluctuation also appears under the ACCEL curriculum (Figure 21), where the agent simply arrives more slowly at its maximum but then exhibits the identical up-and-down pattern.
-In other words, these oscillations reflect an inherent tendency of the RL policy to suffer occasional catastrophic failures in held-out levels, rather than a pathology introduced by our task-prioritization strategy.
-Addressing and damping these fluctuations in the student agent’s performance is an important direction for future work.
-Appendix O Limitations
-As with any research, our approach presents several limitations that highlight opportunities for future investigation:
-  • Co-Learnability In this paper, we proposed Co-Learnability, simple yet effective method.
-Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
-  • Sampling levels We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods.
-However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
-  • Heuristic weighting We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty.
-While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
-  • Level mutation We adopted the same approach as ACCEL which employs random mutations.
-Our approach focused on improving task priority rather than advancing environment parameter mutation mechanisms.
-The current random mutation process creates a clear opportunity for future work to advance these methods using evolutionary algorithms such as CMA-ES [43], so the Editor can offer the agent more informative tasks.
-  • Extension to other RL algorithms Our experiments have been conducted with PPO-based student agents.
-While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum.
-Off-policy methods such as SAC [44] and TD3 [45] feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
-Moreover, Model-based approaches like DreamerV3 [46] build world models for planning, offering complementary insights into curriculum design.
-Evaluating these methods under our curriculum framework remains an important direction for future work.
-  • Extend to other UED methods We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost.
-While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR [6] and CENIE [20]) and conduct comprehensive experiments to assess its performance in these settings.
-  • Evaluation in other domains Our experimental validation has been conducted in MiniGrid [16, 4] and BipedalWalker [23, 7].
-Evaluating TRACED in other domains such as Car Racing [23, 6], physical robotics applications [47], and reasoning task ARC [48] would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
-Appendix P Experimental Setup and Reproducibility
-All experiments were conducted using Python 3.8.20 on Ubuntu 22.04.4 LTS.
-The hardware setup included an AMD EPYC 7543 32-core processor, 80GB of RAM, and an NVIDIA A100 GPU.
-Experiments were run on a shared server with 8 GPUs, with each experiment using a single GPU.
-We implemented all experiments in Python 3.8.20, using PyTorch(v1.9.0+cu111).
-The environment stack was built using MiniGrid (v1.0.1), Gym (v0.15.7).
+Appendix P Numerical Results
+P.1 Minigrid Environment
+Table 11: MiniGrid Zero-Shot Solved Rates. Bold indicates the best result per task; underline indicates the second-best.
+Environment Update DR PLR ADD ACCEL TRACED
+16Rooms 10k 0.4 $\pm$ 0.12 0.81 $\pm$ 0.06 0.72 $\pm$ 0.1 0.76 $\pm$ 0.12 0.53 $\pm$ 0.14
+20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
+16Rooms2 10k 0.1 $\pm$ 0.09 0.42 $\pm$ 0.11 0.14 $\pm$ 0.05 0.48 $\pm$ 0.13 0.58 $\pm$ 0.13
+20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
+SimpleCrossing 10k 0.58 $\pm$ 0.05 0.72 $\pm$ 0.03 0.63 $\pm$ 0.03 0.81 $\pm$ 0.03 0.85 $\pm$ 0.03
+20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
+FourRooms 10k 0.32 $\pm$ 0.04 0.47 $\pm$ 0.01 0.39 $\pm$ 0.03 0.46 $\pm$ 0.02 0.46 $\pm$ 0.02
+20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
+SmallCorridor 10k 0.3 $\pm$ 0.07 0.52 $\pm$ 0.13 0.47 $\pm$ 0.1 0.37 $\pm$ 0.14 0.49 $\pm$ 0.13
+20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
+LargeCorridor 10k 0.36 $\pm$ 0.13 0.33 $\pm$ 0.13 0.35 $\pm$ 0.09 0.44 $\pm$ 0.15 0.43 $\pm$ 0.11
+20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
+Labyrinth 10k 0.1 $\pm$ 0.1 0.5 $\pm$ 0.15 0.22 $\pm$ 0.13 0.6 $\pm$ 0.15 0.93 $\pm$ 0.03
+20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
+Labyrinth2 10k 0.1 $\pm$ 0.1 0.28 $\pm$ 0.13 0.0 $\pm$ 0.0 0.52 $\pm$ 0.14 0.96 $\pm$ 0.03
+20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
+Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
+20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
+Maze2 10k 0.19 $\pm$ 0.12 0.43 $\pm$ 0.14 0.09 $\pm$ 0.08 0.44 $\pm$ 0.15 0.6 $\pm$ 0.1
+20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
+Maze3 10k 0.27 $\pm$ 0.13 0.57 $\pm$ 0.13 0.2 $\pm$ 0.1 0.78 $\pm$ 0.13 0.81 $\pm$ 0.06
+20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
+PerfectMaze(M) 10k 0.06 $\pm$ 0.02 0.24 $\pm$ 0.05 0.17 $\pm$ 0.08 0.41 $\pm$ 0.05 0.64 $\pm$ 0.07
+20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
+Mean 10k 0.23 $\pm$ 0.05 0.45 $\pm$ 0.04 0.28 $\pm$ 0.04 0.54 $\pm$ 0.03 0.66 $\pm$ 0.03
+20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
+Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal et al.
+(2021).
+We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k).
+At 10k, TRACED achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best in one maze (Maze2).
+Averaged across all twelve mazes, TRACED attains a mean solved rate of 0.66, outperforming all baseline configurations.
+Table 12: MiniGrid Ablation Results. Bold indicates the best result per task; underline indicates the second-best.
+Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
+16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
+16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
+SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
+FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
+SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
+LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
+Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
+Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
+Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
+Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
+Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
+PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
+Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
+Table 12 presents zero-shot solved rates on twelve held-out MiniGrid mazes for TRACED and its ablations at 10k updates, averaged over 10 seeds (mean $\pm$ s.e.).
+Full TRACED achieves the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the remaining six, yielding a mean solved rate of 0.70 and outperforming every variant.
+These results indicate that both the transition-prediction error and Co-Learnability are critical to TRACED’s strong, consistent performance across diverse tasks.
+Table 13: PerfectMaze Zero-Shot Solved Rates. Bold indicates the best result per environment; underline indicates the second-best.
+Environment Update DR PLR ADD ACCEL TRACED
+PerfectMaze(Large) 10k 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01 0.04 $\pm$ 0.04 0.09 $\pm$ 0.03 0.27 $\pm$ 0.07
+20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
+PerfectMaze(XL) 10k 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.02 0.01 $\pm$ 0.01 0.1 $\pm$ 0.04
+20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
+Table 13 compares zero‐shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large‐scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL, averaged over 10 seeds (mean $\pm$ s.e.).
+After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third‐highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k.
+These results underscore TRACED’s ability to scale to extreme maze sizes.
+P.2 BipedalWalker Environment
+Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.).
+ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Teoh et al.
+(2024) to our ACCEL implementation.
+At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.
+Table 14: BipedalWalker Zero-Shot Test Returns. Bold indicates the best result per terrain; underline indicates the second-best.
+Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
+Basic 10k 112.56 $\pm$ 62.99 131.63 $\pm$ 56.58 119.43 $\pm$ 63.72 281.65 $\pm$ 5.25 273.56 $\pm$ 1.5 293.67 $\pm$ 3.56
+20k 68.17 $\pm$ 44.71 97.67 $\pm$ 61.73 63.64 $\pm$ 62.71 281.85 $\pm$ 3.72 275.04 $\pm$ 3.19
+Hardcore 10k -19.67 $\pm$ 7.35 -17.36 $\pm$ 5.99 3.83 $\pm$ 9.07 37.59 $\pm$ 15.0 66.83 $\pm$ 17.48 86.83 $\pm$ 17.96
+20k -16.98 $\pm$ 5.8 -23.7 $\pm$ 2.05 10.48 $\pm$ 20.12 59.23 $\pm$ 25.5 84.09 $\pm$ 34.32
+Stairs 10k -17.6 $\pm$ 7.21 -7.23 $\pm$ 7.16 -5.87 $\pm$ 9.07 -38.71 $\pm$ 10.54 -30.4 $\pm$ 8.35 -29.0 $\pm$ 10.4
+20k -9.11 $\pm$ 7.67 -10.2 $\pm$ 4.5 -0.66 $\pm$ 9.2 -46.34 $\pm$ 5.78 -36.39 $\pm$ 11.48
+PitGap 10k -34.39 $\pm$ 15.66 -46.21 $\pm$ 15.69 -17.8 $\pm$ 13.82 -65.07 $\pm$ 7.57 -39.81 $\pm$ 18.75 -39.26 $\pm$ 11.42
+20k -26.03 $\pm$ 10.98 -48.72 $\pm$ 12.23 -7.65 $\pm$ 7.52 -64.89 $\pm$ 18.93 -45.92 $\pm$ 33.9
+Stump 10k -24.99 $\pm$ 6.76 -27.74 $\pm$ 6.33 -26.15 $\pm$ 8.49 -79.18 $\pm$ 5.45 -60.05 $\pm$ 10.96 34.16 $\pm$ 54.58
+20k -25.52 $\pm$ 12.39 -23.61 $\pm$ 2.86 3.26 $\pm$ 19.1 -67.18 $\pm$ 15.56 -46.34 $\pm$ 176.21
+Roughness 10k -0.66 $\pm$ 12.01 3.55 $\pm$ 9.91 20.61 $\pm$ 13.78 161.72 $\pm$ 28.36 174.92 $\pm$ 2.25 193.29 $\pm$ 21.6
+20k 0.19 $\pm$ 8.18 -2.73 $\pm$ 9.51 18.71 $\pm$ 21.59 213.48 $\pm$ 7.69 224.4 $\pm$ 8.63
+Mean 10k 2.54 $\pm$ 15.44 6.1 $\pm$ 10.69 14.4 $\pm$ 14.05 49.67 $\pm$ 11.24 64.18 89.95 $\pm$ 12.95
+20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
+Table 15 reports zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED, averaged over five seeds.
+A trajectory is considered “solved” if its return exceeds 230.
+At 10k updates, TRACED achieves the highest solved rate on five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36, substantially higher than ACCEL’s 0.29 and the other baselines.
+No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
+Table 15: BipedalWalker Zero‐Shot Solved Rates. Bold indicates the best per terrain; underline indicates the second‐best.
+Environment Update DR PLR ADD ACCEL TRACED
+Basic 10k 0.35 $\pm$ 0.2 0.43 $\pm$ 0.21 0.34 $\pm$ 0.21 0.98 $\pm$ 0.02 1.00 $\pm$ 0.0
+20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
+Hardcore 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.01 $\pm$ 0.01 0.15 $\pm$ 0.03 0.28 $\pm$ 0.05
+20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
+Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
+20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
+PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
+20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
+Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
+20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
+Roughness 10k 0.01 $\pm$ 0.01 0.0 $\pm$ 0.01 0.01 $\pm$ 0.01 0.56 $\pm$ 0.1 0.65 $\pm$ 0.08
+20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
+Mean 10k 0.06 $\pm$ 0.04 0.07 $\pm$ 0.04 0.06 $\pm$ 0.04 0.29 $\pm$ 0.02 0.36 $\pm$ 0.03
+20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03

### v4

@@
 1 Introduction
-Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (Sutton et al., 1998).
+Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (sutton1998reinforcement).
 Ideally, we want agents that generalize robustly to a broad range of unseen environments.
-However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (Kirk et al., 2023; Korkmaz, 2024).
-Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (Dennis et al., 2020).
-A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022; Azad et al., 2023b; Mediratta et al., 2023; Erlebach & Cook, 2024).
+However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (rl_generalization; rl_generalization2).
+Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (PAIRED).
+A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (PAIRED; PLR; REPAIRED; ACCEL; azad2023clutr; Stabilizing; erlebach2024raccoon).
 Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains.
-Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (Rutherford et al., 2024).
+Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (rutherford2024no).
 In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics.
@@
 We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others.
-For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (Ramírez et al., 2013; Costa et al., 2000; Lemhöfer & Dijkstra, 2004), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick & and, 2005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability.
+For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (ramirez2013cross; co_ex1; co_ex2), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick15012005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability.
 We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
 We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1).
-We evaluate TRACED on procedurally generated MiniGrid (MG) (Chevalier-Boisvert et al., 2023) and BipedalWalker (BW) (Romac et al., 2021).
-In MG, we compare against Domain Randomization (DR) (Jakobi, 1997), PLR (Jiang et al., 2021b), ADD (Chung et al., 2024), and ACCEL (Parker-Holder et al., 2022) (the strongest baseline).
-In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Teoh et al., 2024).
+We evaluate TRACED on procedurally generated MiniGrid (MG) (MinigridMiniworld23) and BipedalWalker (BW) (BipedalWalker).
+In MG, we compare against Domain Randomization (DR) (DR), PLR (PLR), ADD (Chung2024AdversarialDiffusion), and ACCEL (ACCEL) (the strongest baseline).
+In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Jayden2024CENIE).
 TRACED surpasses all baselines in mean solved rate across 12 MiniGrid mazes, BipedalWalker, and even the extreme PerfectMaze variants.
@@
 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
-We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al.
-(2020).
+We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following PAIRED.
 A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space.
@@
 2.2 Unsupervised Environment Design (UED)
-UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (Dennis et al., 2020).
-Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (Jiang et al., 2021b; Parker-Holder et al., 2022; Teoh et al., 2024).
+UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (PAIRED).
+Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (PLR; ACCEL; Jayden2024CENIE).
 The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
 Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics.
-For example, PLR (Jiang et al., 2021a) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
+For example, PLR (REPAIRED) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
 PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode.
@@
 MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target.
-Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (Jiang et al., 2021b).
+Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (PLR).
 3 TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design
@@
 For each sampled task instance $\tau$, we compute $\widehat{\mathrm{Regret}}(\tau)$ and append it to the task difficulty buffer (TDB), which is subsequently used to drive curriculum updates (Figure 2).
+Detailed theoretical analysis of the relationship between term (iii) in Eq.
+2 and ATPL is provided in Appendix R, where we show that the dynamics-induced component of the future-value gap is upper-bounded by ATPL.
+This establishes ATPL as a principled correction term for regret approximation.
 Figure 2: Task Difficulty Calculation Workflow.
@@
 3.2 Task Priority Construction
-A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes et al., 2024).
+A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes2024OpenEndednessICML).
 Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others.
@@
 This ensures that $\mathrm{TaskDifficulty}(i,k)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
+Note that a difficulty value for task $i$ is well-defined at time $k$ only if the task has been sampled at least once on or before $k$.
+This difficulty estimate becomes available for use when constructing the curriculum at time k+1k\!+\!1.
 Definition 3 (Co‑Learnability).
@@
 A positive Co-Learnability value indicates that visiting $i$ yields transfer benefits.
+Since $\mathrm{CoLearnability}_{i}(k)$ depends on difficulty updates that occur at time $k+1$, it becomes available beginning at timestep $k+2$.
 Definition 4 (Task Priority).
-We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (Parker-Holder et al., 2022):
+We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (ACCEL):
 $\mathrm{TaskPriority}(i,t)=\mathrm{Rank}\Bigl(\mathrm{TaskDifficulty}(i,t)+\beta\,\mathrm{CoLearnability}(i,t)\Bigr)$ (7)
@@
 At each step, we sample tasks inversely to their priority (e.g., $p(i\mid t)\propto 1/\mathrm{TaskPriority}(i,t)$), so lower ranks (higher priority) are selected more often.
-The rank transform prevents outliers from dominating the sampling distribution.
+The rank transform mitigates the influence of outliers by discarding the absolute magnitude of raw scores and retaining only their relative ordering.
+Without this transformation, a single task with an anomalously large difficulty or Co-Learnability value can dominate the sampling distribution, collapsing the curriculum toward that task.
 3.3 Overall UED Workflow
@@
 (3) Task Mutation: the lowest‑priority levels are mutated into new variants and reinserted into the buffer.
-The overall UED algorithm follows the ACCEL loop (Parker-Holder et al., 2022), with the sole change that task scoring uses our Task Priority (Eq.
+The overall UED algorithm follows the ACCEL loop (ACCEL), with the sole change that task scoring uses our Task Priority (Eq.
 7) in place of Positive Value Loss (PVL) alone (Figure 3).
@@
 In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD.
-In BW only, we additionally include the recent sota method CENIE (Teoh et al., 2024).
-We exclude SFL (Rutherford et al., 2024) because, on MiniGrid, it does not outperform PLR (Rutherford et al., 2024).
+In BW only, we additionally include the recent sota method CENIE (Jayden2024CENIE).
+While SFL (rutherford2024no) is not included in the main experimental suite, we provide a comprehensive comparison against SFL in the multi-agent JaxNav environment in Appendix Q, following its original evaluation protocol.
 Detailed descriptions of all baselines can be found in Appendix K.
 For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels.
-To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal et al., 2021).
+To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal2021Reliable).
 The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible.
 Accordingly, higher IQM and lower optimality gap values indicate better performance.
-All methods use Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the student agent.
-Following prior work (Parker-Holder et al., 2022; Teoh et al., 2024), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F.
+Shaded regions in all plots denote 95% confidence intervals.
+All methods use Proximal Policy Optimization (PPO) (Schulman2017PPO) as the student agent.
+Following prior work (ACCEL; Jayden2024CENIE), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F.
 Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16.
@@
 4.1 Partially Observable Navigation
-We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (Chevalier-Boisvert et al., 2023), as in prior UED work.
+We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (MinigridMiniworld23), as in prior UED work.
 Four representatives are shown in Figure 5.
@@
 On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes.
-Detailed per-baseline results for both benchmarks are provided in Appendix P.
+Complete per-baseline numerical results for both benchmarks are reported in Appendix P, and the full set of aggregate metrics (Median, IQM, Mean, and Optimality Gap) for PerfectMaze is shown in Figure 22.
 4.2 Walking in Challenging Terrain
-We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman et al., 2016), as modified by Wang et al.
-(2019).
+We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman2016Gym), as modified by POET.
 This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty.
@@
 To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds.
-Results are reported following Agarwal et al.
-(2021).
+Results are reported following Agarwal2021Reliable.
 On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance.
@@
 To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters.
-Following Teoh et al.
-(2024), we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5).
+Following Jayden2024CENIE, we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5).
 Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging.
@@
 We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel.
-The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively.
+The only difference between our experimental setup and ACCEL (ACCEL) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively.
 To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers.
@@
 Algorithm 1 summarizes our overall UED procedure under the TRACED framework.
-We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority.
+We followed ACCEL (ACCEL) in all respects except the procedure used to compute task priority.
 At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
@@
 In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths.
-SimpleCrossing and FourRooms are adapted from MiniGrid (Chevalier-Boisvert et al., 2023), and PerfectMaze from REPAIRED (Jiang et al., 2021a).
-The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (Chevalier-Boisvert et al., 2023).
+SimpleCrossing and FourRooms are adapted from MiniGrid (MinigridMiniworld23), and PerfectMaze from REPAIRED (REPAIRED).
+The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (MinigridMiniworld23).
 Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself.
@@
 If the agent fails to reach the goal within the time limit, it does not receive a reward.
-To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (Dennis et al., 2020; Jiang et al., 2021a).
+To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (PAIRED; REPAIRED).
 Figure 20 displays all evaluation tasks.
 H.2 BipedalWalker Environment
-The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman et al., 2016).
+The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman2016Gym).
 At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead.
@@
 and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output.
-We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (Ha & Schmidhuber, 2018; Oh et al., 2015).
+We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (worldmodels; ohjunhyuk2015action).
 For walking in challenging terrain (BipedalWalker), the model instead concatenates the raw state vector $s_{t}$ with $a_{t}$ and feeds this into an LSTM (hidden size 128, 1 layer); the recurrent output is then linearly mapped to the predicted next state $\hat{s}_{t+1}$.
@@
 $\ell_{t}=\begin{cases}\|\,\hat{s}_{t+1}-s_{t+1}\|_{1},&\text{in MiniGrid environments},\\ \|\hat{s}_{t+1}-s_{t+1}\|_{2}^{2},&\text{in BipedalWalker}.\end{cases}$
-We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (Zhao et al., 2016), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (Mnih et al., 2015).
-We adopt ACCEL’s network architecture (Parker-Holder et al., 2022), differing only in the LSTM design choice.
+We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (zhao2016loss), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (mnih2015human).
+We adopt ACCEL’s network architecture (ACCEL), differing only in the LSTM design choice.
 The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment.
@@
 Appendix J Related Works
-Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (Wang et al., 2019; 2020; Dennis et al., 2020; Chung et al., 2024).
-PAIRED (Dennis et al., 2020) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
+Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (POET; EnhancedPOET; PAIRED; Chung2024AdversarialDiffusion).
+PAIRED (PAIRED) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
 However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula.
-To mitigate these limitations, Mediratta et al.
-(2023) introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (Azad et al., 2023a) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
-Another line of work is based on Prioritized Level Replay (PLR) (Jiang et al., 2021b), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
+To mitigate these limitations, Stabilizing introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (CLUTR) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
+Another line of work is based on Prioritized Level Replay (PLR) (PLR), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
 This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent.
-PLR  (Jiang et al., 2021a) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
+PLR  (REPAIRED) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
 This leads to better performance on unseen environments counterintuitively.
-ACCEL (Parker-Holder et al., 2022) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
+ACCEL (ACCEL) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
 This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks.
-However, these methods (Jiang et al., 2021b; a; Parker-Holder et al., 2022) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
-Moving beyond regret-based approaches, CENIE (Teoh et al., 2024) integrates novelty with regret to expose agents to more diverse learning situations.
+However, these methods (PLR; REPAIRED; ACCEL) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
+Moving beyond regret-based approaches, CENIE (Jayden2024CENIE) integrates novelty with regret to expose agents to more diverse learning situations.
 Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer.
 This curriculum-awareness leads to better exploration and generalization.
-DIPLR (Li et al., 2023) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories.
-MBeDED (Li et al., 2024) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
+DIPLR (DIPLR) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories.
+MBeDED (Marginal) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
 It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task.
+SFL (rutherford2024no) assigns a learnability score to each level, computed from the empirical success rate achieved by evaluating the agent on sampled levels.
+Levels whose success rates fall in the intermediate range, neither too easy nor too difficult, are prioritized, guiding the agent toward tasks that are most conducive to further learning.
+A related idea appears in GoalGAN (Florensa2018GoalGAN), which is not a UED method but similarly adapts task difficulty by training a generative model to propose goals of suitable challenge.
 So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
 UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains.
-For Multi-Agent Reinforcement Learning, MAESTRO (Samvelyan et al., 2023) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
-In contrast, RACCOON (Erlebach & Cook, 2024) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
-For meta-RL, GROOVE (Jackson et al., 2023) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
-DIVA (Costales & Nikolaidis, 2024) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
-In sim-to-real settings, GCL (Wang et al., 2024) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
-Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (Narvekar et al., 2020).
-Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (Klink et al., 2020).
-Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (Eimer et al., 2021; Klink et al., 2022).
-WAKER (Rigter et al., 2024) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
+For Multi-Agent Reinforcement Learning, MAESTRO (samvelyan2023maestro) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
+In contrast, RACCOON (erlebach2024raccoon) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
+For meta-RL, GROOVE (groove) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
+DIVA (DIVA) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
+In sim-to-real settings, GCL (wang2024grounded) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
+Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (CurriculmRLSurvey).
+Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (SPDRL).
+Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (SPaCE; currot).
+WAKER (rigter2024rewardfree) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
+Akin to the concept underlying SFL (rutherford2024no), ProCuRL (ProximalCurriculum2023) operationalizes the pedagogical notion of the Zone of Proximal Development by using probability of success score-based scoring combined with softmax sampling to prioritize tasks of intermediate difficulty.
 However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
@@
 The agent is trained on these randomly sampled tasks.
-Robust PLR (PLR(Jiang et al., 2021a) was introduced as an extension of the original PLR (Jiang et al., 2021b).
+Robust PLR (PLR(REPAIRED) was introduced as an extension of the original PLR (PLR).
 While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay.
@@
 By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
-Adversarially Compounding Complexity by Editing Levels (ACCEL) (Parker-Holder et al., 2022) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
+Adversarially Compounding Complexity by Editing Levels (ACCEL) (ACCEL) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
 This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret.
@@
 Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
-Coverage-based Evaluation of Novelty In Environment (CENIE) (Teoh et al., 2024) introduces environment novelty as a complementary objective to regret in UED.
+Coverage-based Evaluation of Novelty In Environment (CENIE) (Jayden2024CENIE) introduces environment novelty as a complementary objective to regret in UED.
 CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum.
@@
 CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
-Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung et al., 2024) frames environment generation as a diffusion process steered by the agent’s regret.
+Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung2024AdversarialDiffusion) frames environment generation as a diffusion process steered by the agent’s regret.
 Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
-In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (Jiang et al., 2022), and assessed ADD with the ADD codebase (Chung et al., 2024).
+In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (dcd_github), and assessed ADD with the ADD codebase (Chung2024AdversarialDiffusion).
 CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
@@
 While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum.
-Off-policy methods such as SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
-Moreover, Model-based approaches like DreamerV3 (Hafner et al., 2025) build world models for planning, offering complementary insights into curriculum design.
+Off-policy methods such as SAC (haarnoja2018soft) and TD3 (fujimoto2018addressing) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
+Moreover, Model-based approaches like DreamerV3 (hafner2025) build world models for planning, offering complementary insights into curriculum design.
 Evaluating these methods under our curriculum framework remains an important direction for future work.
@@
 We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost.
-While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (Jiang et al., 2021a) and CENIE (Teoh et al., 2024)) and conduct comprehensive experiments to assess its performance in these settings.
+While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (REPAIRED) and CENIE (Jayden2024CENIE)) and conduct comprehensive experiments to assess its performance in these settings.
 Evaluation in other domains.
-Our experimental validation has been conducted in MiniGrid (Chevalier-Boisvert et al., 2023; Dennis et al., 2020) and BipedalWalker (Brockman et al., 2016; Parker-Holder et al., 2022).
-Evaluating TRACED in other domains such as Car Racing (Brockman et al., 2016; Jiang et al., 2021a), physical robotics applications (Mahmood et al., 2018), and reasoning task ARC (Chollet, 2019) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
+Our experimental validation has been conducted in MiniGrid (MinigridMiniworld23; PAIRED) and BipedalWalker (Brockman2016Gym; ACCEL).
+Evaluating TRACED in other domains such as Car Racing (Brockman2016Gym; REPAIRED), physical robotics applications (mahmood2018benchmarking), and reasoning task ARC (chollet2019measure) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
 Appendix M Experimental Setup and Reproducibility
@@
 Table 10 summarizes the hyperparameters used.
-Unless otherwise noted, values are adopted directly from the DCD repository (Jiang et al., 2022).
+Unless otherwise noted, values are adopted directly from the DCD repository (dcd_github).
 The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4).
@@
 20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
-Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal et al.
-(2021).
+Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal2021Reliable.
 We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k).
@@
 These results underscore TRACED’s ability to scale to extreme maze sizes.
+(a) PerfectMazeXL
+(b) PerfectMazeLarge
+Figure 22: Aggregate zero-shot performance on PerfectMaze. Full metric curves (Median, IQM, Mean, Optimality Gap) for PerfectMazeXL (top) and PerfectMazeLarge (bottom).
+Figure 22 presents the full aggregate metrics for both PerfectMaze benchmarks.
+Despite the high variance inherent to these large-scale tasks, TRACED achieves the strongest 10k performance and closely matches the best 20k performance among all baselines.
 P.2 BipedalWalker Environment
 Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.).
-ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Teoh et al.
-(2024) to our ACCEL implementation.
+ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Jayden2024CENIE to our ACCEL implementation.
 At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.
@@
 20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03
+Appendix Q Comparison with SFL
+(a) CVaR across $\alpha$ levels.
+(b) Mean win rate on the evaluation set.
+Figure 23: TRACED vs. SFL and Robust PLR on multi-agent JaxNav. TRACED consistently outperforms Robust PLR and achieves performance comparable to SFL across all metrics.
+Sampling for Learnability (SFL) (rutherford2024no) is a learnability-driven UED method that prioritizes levels with intermediate difficulty by selecting tasks with high learnability scores, computed as p(1p)p(1-p) where $p$ is the agent’s empirical success rate on a given level obtained from rollouts.
+We evaluate TRACED on the multi-agent JaxNav benchmark following the evaluation protocol introduced in rutherford2024no, including CVaR and mean win rate computed over sampled evaluation levels.
+While the original SFL experiments trained baselines for 22,850 PPO updates, we perform our comparison using 4,550 updates for computational efficiency.
+This training regime corresponds to a phase in which SFL is reported to exhibit strong performance relative to other UED baselines (rutherford2024no).
+Aside from the inclusion of TRACED, all training and evaluation details (e.g., level sampling, CVaR estimation, and construction of evaluation sets) match those in rutherford2024no.
+Results are averaged over three seeds.
+Figure 24 shows CVaR across several risk levels and the mean win rate on the evaluation set.
+In addition to SFL, we include PLR (Robust PLR), the strongest baseline used in rutherford2024no.
+TRACED consistently outperforms Robust PLR across all CVaR levels and achieves performance comparable to SFL in both CVaR and mean win rate.
+Notably, TRACED attains this performance without relying on any learnability-driven evaluation during training.
+These results suggest that TRACED captures many of the benefits of learnability-based curricula while depending solely on a regret-based difficulty signal.
+Incorporating SFL-style evaluation tools, such as those used to quantify Co-Learnability, may further enhance TRACED’s performance in complex multi-agent settings.
+Appendix R Theoretical Analysis of ATPL
+We restate the one-step regret decomposition (Eq. 2 in the main paper):
+$\mathrm{Regret}(s,a)$ $=\underbrace{V^{*}(s)-\hat{V}^{*}(s)}_{\text{(i) Value estimation error}}+\underbrace{r(s,a^{*})-r(s,a)}_{\text{(ii) Reward gap}}$
+$\quad+\gamma\,\underbrace{\Bigl(\mathbb{E}_{s^{\prime\prime}\sim\hat{P}(\cdot\mid s,a^{*})}\!\bigl[\hat{V}^{*}(s^{\prime\prime})\bigr]-\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\!\bigl[V^{\pi}(s^{\prime})\bigr]\Bigr)}_{\text{(iii) Future value gap}}\,$,
+Our objective is to isolate and control the portion of term (iii) that arises specifically from the mismatch between the true transition kernel $P$ and the learned transition model $\hat{P}$.
+In particular, we focus on the discrepancy between the value predictions under $P$ and $\hat{P}$, and show that this dynamics-induced error can be bounded by the transition loss used in ATPL.
+R.1 Assumptions
+Assumption 1 (Lipschitz value function).
+Let $(\mathcal{S},d)$ be the state space equipped with a metric $d(\cdot,\cdot)$.
+The approximate optimal value function $\widehat{V}^{*}:\mathcal{S}\to\mathbb{R}$ is $L_{V}$-Lipschitz with respect to $d$, i.e.,
+$\bigl|\widehat{V}^{*}(s_{1})-\widehat{V}^{*}(s_{2})\bigr|\leq L_{V}\,d(s_{1},s_{2}),\qquad\forall s_{1},s_{2}\in\mathcal{S}.$
+Assumption 2 (Loss dominating the metric).
+Let $\ell:\mathcal{S}\times\mathcal{S}\to[0,\infty)$ be the reconstruction loss used to train the transition model (e.g., $\ell(x,y)=\|x-y\|^{2}$).
+We assume that there exists a constant $C_{\ell}>0$ such that
+$d(x,y)\leq C_{\ell}\,\ell(x,y),\qquad\forall x,y\in\mathcal{S}.$
+Definition 5 (Coupling-based transition loss).
+For each state-action pair $(s,a)$, let $P(\cdot\mid s,a)$ and $\hat{P}(\cdot\mid s,a)$ denote the true and learned transition kernels, respectively.
+A coupling of $P(\cdot\mid s,a)$ and $\hat{P}(\cdot\mid s,a)$ is any joint distribution $\Gamma_{s,a}$ on $\mathcal{S}\times\mathcal{S}$ such that its marginals satisfy
+$(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}\quad\Longrightarrow\quad s^{+}\sim P(\cdot\mid s,a),\qquad\hat{s}^{+}\sim\hat{P}(\cdot\mid s,a).$
+Given such a coupling $\Gamma_{s,a}$, we define the one-step transition loss as
+$L_{\mathrm{trans}}(s,a):=\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\ell(s^{+},\hat{s}^{+})\bigr].$
+R.2 Bounding the Dynamics Mismatch Term
+Lemma 1.
+Define the dynamics mismatch term
+$\Delta_{\mathrm{dyn}}(s,a):=\Bigl|\mathbb{E}_{\hat{P}(\cdot\mid s,a)}\bigl[\widehat{V}^{*}(s^{\prime\prime})\bigr]-\mathbb{E}_{P(\cdot\mid s,a)}\bigl[\widehat{V}^{*}(s^{\prime})\bigr]\Bigr|.$
+Under Assumptions 1 and 2, for any coupling $\Gamma_{s,a}$ of $P(\cdot\mid s,a)$ and $\hat{P}(\cdot\mid s,a)$, we have
+$\Delta_{\mathrm{dyn}}(s,a)\;\leq\;L_{V}\,C_{\ell}\,L_{\mathrm{trans}}(s,a).$
+Proof.
+Let $(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}$ be a coupling of $P(\cdot\mid s,a)$ and $\hat{P}(\cdot\mid s,a)$, i.e., $s^{+}\sim P(\cdot\mid s,a)$ and $\hat{s}^{+}\sim\hat{P}(\cdot\mid s,a)$.
+Then, by the definition of the marginal distributions,
+$\mathbb{E}_{\hat{P}(\cdot\mid s,a)}\bigl[\widehat{V}^{*}(s^{\prime\prime})\bigr]=\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\widehat{V}^{*}(\hat{s}^{+})\bigr],$
+and
+$\mathbb{E}_{P(\cdot\mid s,a)}\bigl[\widehat{V}^{*}(s^{\prime})\bigr]=\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\widehat{V}^{*}(s^{+})\bigr].$
+Hence
+$\Delta_{\mathrm{dyn}}(s,a)$ $=\Bigl|\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\widehat{V}^{*}(\hat{s}^{+})\bigr]-\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\widehat{V}^{*}(s^{+})\bigr]\Bigr|$
+$=\Bigl|\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\widehat{V}^{*}(\hat{s}^{+})-\widehat{V}^{*}(s^{+})\bigr]\Bigr|$
+$\leq\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\Bigl[\bigl|\widehat{V}^{*}(\hat{s}^{+})-\widehat{V}^{*}(s^{+})\bigr|\Bigr]$
+$\leq L_{V}\,\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[d(s^{+},\hat{s}^{+})\bigr]\qquad\text{(by Lipschitzness of $\widehat{V}^{*}$)}$
+$\leq L_{V}\,C_{\ell}\,\mathbb{E}_{(s^{+},\hat{s}^{+})\sim\Gamma_{s,a}}\bigl[\ell(s^{+},\hat{s}^{+})\bigr]\qquad\text{(by the dominance $d\leq C_{\ell}\ell$)}$
+$=L_{V}\,C_{\ell}\,L_{\mathrm{trans}}(s,a).$
+R.3 Regret Approximation
+We compare two regret proxies:
+$\widehat{\mathrm{Regret}}_{\mathrm{PVL}}(\tau):=\mathrm{PVL}(\tau),\qquad\widehat{\mathrm{Regret}}_{\mathrm{ATPL}}(\tau):=\mathrm{PVL}(\tau)+\alpha\,\mathrm{ATPL}(\tau).$
+Theorem 1 (Regret Approximation Improvement with ATPL).
+For any trajectory $\tau$,
+$\bigl|\mathrm{Regret}(\tau)-\widehat{\mathrm{Regret}}_{\mathrm{PVL}}(\tau)\bigr|\;\leq\;C_{0}+L_{V}C_{\ell}\,\mathrm{ATPL}(\tau),$
+where $C_{0}$ collects the remaining terms in the regret decomposition that do not depend on the transition mismatch. Furthermore,
+$\Bigl|\mathrm{Regret}(\tau)-\widehat{\mathrm{Regret}}_{\mathrm{ATPL}}(\tau)\Bigr|\;\leq\;C_{0}+\bigl|L_{V}C_{\ell}-\alpha\bigr|\,\mathrm{ATPL}(\tau).$
+In particular, choosing $\alpha=L_{V}C_{\ell}$ yields a strictly tighter worst-case approximation bound than PVL alone.
+Proof.
+Lemma 1 bounds the dynamics-induced component of the regret by $L_{V}C_{\ell}\,\mathrm{ATPL}(\tau)$.
+Substituting this bound into the regret decomposition and subtracting $\alpha\,\mathrm{ATPL}(\tau)$ gives the inequality.
+∎
+Interpretation.
+PVL captures only the value-estimation component of regret and assigns zero weight to the error arising from transition-model mismatch.
+Lemma 1 shows that this missing component is controlled by ATPL.
+Theorem 1 then establishes that augmenting PVL with $\alpha\,\mathrm{ATPL}$ reduces the worst-case approximation whenever $\alpha$ is chosen to match the scale of the transition mismatch (i.e., $L_{V}C_{\ell}$).
+Thus, ATPL provides a principled correction for the future-value error, yielding a more faithful regret proxy for unsupervised environment design.
+Appendix S Temporal Consistency of Co-Learnability
+Figure 24: Temporal consistency of Co-Learnability.
+Shadow testing produces aligned Co-Learnability sequences, enabling direct comparison of $\mathrm{CoLearnability}_{i}(t)$ and $\mathrm{CoLearnability}_{i}(t\!+\!1)$.
+The two signals exhibit positive correlation throughout training.
+To examine whether training on tasks with high Co-Learnability at time $t$ continues to facilitate progress at time t+1t\!+\!1, we quantify the temporal consistency of Co-Learnability values.
+A key challenge is that Co-Learnability is only computed for tasks selected by the curriculum at each step, resulting in irregularly sampled time series in which many tasks are not evaluated at both $t$ and t+1t\!+\!1.
+This prevents direct computation of temporal correlation.
+To obtain time-aligned Co-Learnability estimates, we apply a shadow-testing procedure.
+Let $L_{t}$ denote the indices of levels selected by the curriculum at timestep $t$.
+After the update on $L_{t}$, we clone the agent’s parameters to construct a shadow agent.
+While the main learner proceeds normally (so that $L_{t+1}$ may differ from $L_{t}$), the shadow agent is forced to replay the same levels by setting $L_{t+1}=L_{t}$.
+At the next timestep, the shadow agent resumes standard TRACED updates without further intervention.
+This procedure yields aligned pairs of Co-Learnability values, $\mathrm{CoLearnability}_{i}(t)$ and $\mathrm{CoLearnability}_{i}(t+1)$, for the same set of levels across all timesteps, enabling direct temporal comparison.
+Figure 24 shows the resulting time series of Co-Learnability at steps $t$ and t+1t\!+\!1, averaged over three random seeds on multi-agent JaxNav.
+Although both signals exhibit substantial variance early in training, they converge toward small positive values as the policy improves and performance becomes more uniform across levels.
+The temporal correlation between the two estimates is summarized below:
+  • Pearson correlation: $0.3998$ ($p\approx 3.9\times 10^{-87}$)
+  • Spearman correlation: $0.2013$ ($p\approx 5.4\times 10^{-22}$)
+  • Time-averaged Co-Learnability at step $t$: $0.0547$
+  • Time-averaged Co-Learnability at step t+1t\!+\!1: $0.0473$
+The slightly lower mean at step t+1t\!+\!1 is expected, as Co-Learnability values naturally decrease over training due to overall improvement across levels.
+The positive correlations indicate that tasks exhibiting higher Co-Learnability at time $t$ tend to retain elevated Co-Learnability at t+1t\!+\!1, suggesting mild but non-negligible temporal persistency.

### v5

@@
 1 Introduction
-Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (sutton1998reinforcement).
+Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics (Sutton et al., 1998).
 Ideally, we want agents that generalize robustly to a broad range of unseen environments.
-However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (rl_generalization; rl_generalization2).
-Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (PAIRED).
-A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (PAIRED; PLR; REPAIRED; ACCEL; azad2023clutr; Stabilizing; erlebach2024raccoon).
+However, hand‑crafting a training distribution that captures all real‑world variability is intractable, and agents often overfit even large training sets, performing poorly out‑of‑distribution (Kirk et al., 2023; Korkmaz, 2024).
+Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent (Dennis et al., 2020).
+A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design (Dennis et al., 2020; Jiang et al., 2021b; a; Parker-Holder et al., 2022; Azad et al., 2023a; Mediratta et al., 2023; Erlebach and Cook, 2024).
 Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains.
-Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (rutherford2024no).
+Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) (Rutherford et al., 2024).
 In this paper, we refine regret estimation by augmenting PVL with a transition-prediction error term that captures how poorly a learned model predicts the environment’s dynamics.
@@
 We further introduce a metric called Co‑Learnability to quantify how training on one task benefits others.
-For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (ramirez2013cross; co_ex1; co_ex2), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick15012005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability.
+For instance, consider three 100‑word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy (Ramírez et al., 2013; Costa et al., 2000; Lemhöfer and Dijkstra, 2004), reflecting high Co‑Learnability; in contrast, Japanese is typologically distant (Chiswick and Miller, 2005), so gains from Japanese may transfer unefficiently to English, reflecting low Co‑Learnability.
 We present a lightweight estimator of Co‑Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
@@
 We propose TRACED (Transition-aware Regret Approximation with Co-Learnability for Environment Design), which combines refined regret and Co-Learnability to yield a task-priority landscape (Figure 1).
-We evaluate TRACED on procedurally generated MiniGrid (MG) (MinigridMiniworld23) and BipedalWalker (BW) (BipedalWalker).
-In MG, we compare against Domain Randomization (DR) (DR), PLR (PLR), ADD (Chung2024AdversarialDiffusion), and ACCEL (ACCEL) (the strongest baseline).
-In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Jayden2024CENIE).
+We evaluate TRACED on procedurally generated MiniGrid (MG) (Chevalier-Boisvert et al., 2023) and BipedalWalker (BW) (Romac et al., 2021).
+In MG, we compare against Domain Randomization (DR) (Jakobi, 1997), PLR (Jiang et al., 2021b), ADD (Chung et al., 2024), and ACCEL (Parker-Holder et al., 2022) (the strongest baseline).
+In BW, we additionally include the state-of-the-art (SOTA) method CENIE (Teoh et al., 2024).
 TRACED surpasses all baselines in mean solved rate across 12 MiniGrid mazes, BipedalWalker, and even the extreme PerfectMaze variants.
@@
 2.1 Underspecified Partially Observable MDPs (UPOMDPs)
-We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following PAIRED.
+We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al.
+(2020).
 A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space.
@@
 2.2 Unsupervised Environment Design (UED)
-UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (PAIRED).
-Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (PLR; ACCEL; Jayden2024CENIE).
+UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ (Dennis et al., 2020).
+Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning (Jiang et al., 2021b; Parker-Holder et al., 2022; Teoh et al., 2024).
 The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
 Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics.
-For example, PLR (REPAIRED) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
+For example, PLR (Jiang et al., 2021a) evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC).
 PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)‐based Temporal Difference (TD) errors over an episode.
@@
 MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target.
-Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (PLR).
+Other criteria such as policy entropy, one‑step TD error, GAE, policy min‑margin, and policy least‑confidence have also been evaluated in curriculum learning contexts (Jiang et al., 2021b).
 3 TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design
@@
 3.2 Task Priority Construction
-A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes2024OpenEndednessICML).
+A central challenge in UED is to decide which tasks to present to the agent at each step (Hughes et al., 2024).
 Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others.
@@
 Definition 4 (Task Priority).
-We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (ACCEL):
+We combine difficulty and Co-Learnability into a scalar score and then apply a rank transform, which replaces raw values with their relative order (e.g., the largest value receives rank 1, the next largest rank 2, etc.), as in ACCEL (Parker-Holder et al., 2022):
 $\mathrm{TaskPriority}(i,t)=\mathrm{Rank}\Bigl(\mathrm{TaskDifficulty}(i,t)+\beta\,\mathrm{CoLearnability}(i,t)\Bigr)$ (7)
@@
 (3) Task Mutation: the lowest‑priority levels are mutated into new variants and reinserted into the buffer.
-The overall UED algorithm follows the ACCEL loop (ACCEL), with the sole change that task scoring uses our Task Priority (Eq.
+The overall UED algorithm follows the ACCEL loop (Parker-Holder et al., 2022), with the sole change that task scoring uses our Task Priority (Eq.
 7) in place of Positive Value Loss (PVL) alone (Figure 3).
@@
 In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD.
-In BW only, we additionally include the recent sota method CENIE (Jayden2024CENIE).
-While SFL (rutherford2024no) is not included in the main experimental suite, we provide a comprehensive comparison against SFL in the multi-agent JaxNav environment in Appendix Q, following its original evaluation protocol.
+In BW only, we additionally include the recent sota method CENIE (Teoh et al., 2024).
+While SFL (Rutherford et al., 2024) is not included in the main experimental suite, we provide a comprehensive comparison against SFL in the multi-agent JaxNav environment in Appendix Q, following its original evaluation protocol.
 Detailed descriptions of all baselines can be found in Appendix K.
 For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels.
-To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal2021Reliable).
+To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal et al., 2021).
 The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible.
@@
 Shaded regions in all plots denote 95% confidence intervals.
-All methods use Proximal Policy Optimization (PPO) (Schulman2017PPO) as the student agent.
-Following prior work (ACCEL; Jayden2024CENIE), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F.
+All methods use Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the student agent.
+Following prior work (Parker-Holder et al., 2022; Teoh et al., 2024), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F.
 Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16.
@@
 4.1 Partially Observable Navigation
-We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (MinigridMiniworld23), as in prior UED work.
+We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (Chevalier-Boisvert et al., 2023), as in prior UED work.
 Four representatives are shown in Figure 5.
@@
 4.2 Walking in Challenging Terrain
-We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman2016Gym), as modified by POET.
+We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman et al., 2016), as modified by Wang et al.
+(2019).
 This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty.
@@
 To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds.
-Results are reported following Agarwal2021Reliable.
+Results are reported following Agarwal et al.
+(2021).
 On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance.
@@
 To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters.
-Following Jayden2024CENIE, we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5).
+Following Teoh et al.
+(2024), we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5).
 Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging.
@@
 We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret‐guided exploration in open‑ended learning environments.
+Acknowledgments
+This work was supported by IITP (RS-2024-00445087; 10%, RS-2025-25410841; 10%, No.
+2019-0-01842; 10%), NRF (RS-2024-00451162; 20%, RS-2024-00454000; 20%, RS-2025-25419920; 20%), GIST (KH0870; 10%) funded by the Ministry of Science and ICT, Korea.
+Computing resource were supported by GIST SCENT-AI, AICA, KISTI, and NIPA.
 Appendix A Additional Experiments
@@
 We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel.
-The only difference between our experimental setup and ACCEL (ACCEL) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively.
+The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively.
 To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers.
@@
 Algorithm 1 summarizes our overall UED procedure under the TRACED framework.
-We followed ACCEL (ACCEL) in all respects except the procedure used to compute task priority.
+We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority.
 At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
@@
 In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths.
-SimpleCrossing and FourRooms are adapted from MiniGrid (MinigridMiniworld23), and PerfectMaze from REPAIRED (REPAIRED).
-The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (MinigridMiniworld23).
+SimpleCrossing and FourRooms are adapted from MiniGrid (Chevalier-Boisvert et al., 2023), and PerfectMaze from REPAIRED (Jiang et al., 2021a).
+The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (Chevalier-Boisvert et al., 2023).
 Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself.
@@
 If the agent fails to reach the goal within the time limit, it does not receive a reward.
-To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (PAIRED; REPAIRED).
+To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (Dennis et al., 2020; Jiang et al., 2021a).
 Figure 20 displays all evaluation tasks.
 H.2 BipedalWalker Environment
-The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman2016Gym).
+The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman et al., 2016).
 At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead.
@@
 and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output.
-We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (worldmodels; ohjunhyuk2015action).
+We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (Ha and Schmidhuber, 2018; Oh et al., 2015).
 For walking in challenging terrain (BipedalWalker), the model instead concatenates the raw state vector $s_{t}$ with $a_{t}$ and feeds this into an LSTM (hidden size 128, 1 layer); the recurrent output is then linearly mapped to the predicted next state $\hat{s}_{t+1}$.
@@
 $\ell_{t}=\begin{cases}\|\,\hat{s}_{t+1}-s_{t+1}\|_{1},&\text{in MiniGrid environments},\\ \|\hat{s}_{t+1}-s_{t+1}\|_{2}^{2},&\text{in BipedalWalker}.\end{cases}$
-We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (zhao2016loss), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (mnih2015human).
-We adopt ACCEL’s network architecture (ACCEL), differing only in the LSTM design choice.
+We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (Zhao et al., 2016), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (Mnih et al., 2015).
+We adopt ACCEL’s network architecture (Parker-Holder et al., 2022), differing only in the LSTM design choice.
 The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment.
@@
 Appendix J Related Works
-Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (POET; EnhancedPOET; PAIRED; Chung2024AdversarialDiffusion).
-PAIRED (PAIRED) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
+Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (Wang et al., 2019; 2020; Dennis et al., 2020; Chung et al., 2024).
+PAIRED (Dennis et al., 2020) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments.
 However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula.
-To mitigate these limitations, Stabilizing introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (CLUTR) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
-Another line of work is based on Prioritized Level Replay (PLR) (PLR), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
+To mitigate these limitations, Mediratta et al.
+(2023) introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (Azad et al., 2023b) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
+Another line of work is based on Prioritized Level Replay (PLR) (Jiang et al., 2021b), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks.
 This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent.
-PLR  (REPAIRED) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
+PLR  (Jiang et al., 2021a) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer.
 This leads to better performance on unseen environments counterintuitively.
-ACCEL (ACCEL) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
+ACCEL (Parker-Holder et al., 2022) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer.
 This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks.
-However, these methods (PLR; REPAIRED; ACCEL) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
-Moving beyond regret-based approaches, CENIE (Jayden2024CENIE) integrates novelty with regret to expose agents to more diverse learning situations.
+However, these methods (Jiang et al., 2021b; a; Parker-Holder et al., 2022) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
+Moving beyond regret-based approaches, CENIE (Teoh et al., 2024) integrates novelty with regret to expose agents to more diverse learning situations.
 Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer.
 This curriculum-awareness leads to better exploration and generalization.
-DIPLR (DIPLR) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories.
-MBeDED (Marginal) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
+DIPLR (Li et al., 2023) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories.
+MBeDED (Li et al., 2024) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task.
 It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task.
-SFL (rutherford2024no) assigns a learnability score to each level, computed from the empirical success rate achieved by evaluating the agent on sampled levels.
+SFL (Rutherford et al., 2024) assigns a learnability score to each level, computed from the empirical success rate achieved by evaluating the agent on sampled levels.
 Levels whose success rates fall in the intermediate range, neither too easy nor too difficult, are prioritized, guiding the agent toward tasks that are most conducive to further learning.
-A related idea appears in GoalGAN (Florensa2018GoalGAN), which is not a UED method but similarly adapts task difficulty by training a generative model to propose goals of suitable challenge.
+A related idea appears in GoalGAN (Florensa et al., 2018), which is not a UED method but similarly adapts task difficulty by training a generative model to propose goals of suitable challenge.
 So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
 UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains.
-For Multi-Agent Reinforcement Learning, MAESTRO (samvelyan2023maestro) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
-In contrast, RACCOON (erlebach2024raccoon) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
-For meta-RL, GROOVE (groove) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
-DIVA (DIVA) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
-In sim-to-real settings, GCL (wang2024grounded) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
-Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (CurriculmRLSurvey).
-Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (SPDRL).
-Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (SPaCE; currot).
-WAKER (rigter2024rewardfree) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
-Akin to the concept underlying SFL (rutherford2024no), ProCuRL (ProximalCurriculum2023) operationalizes the pedagogical notion of the Zone of Proximal Development by using probability of success score-based scoring combined with softmax sampling to prioritize tasks of intermediate difficulty.
+For Multi-Agent Reinforcement Learning, MAESTRO (Samvelyan et al., 2023) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations.
+In contrast, RACCOON (Erlebach and Cook, 2024) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration.
+For meta-RL, GROOVE (Jackson et al., 2023) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training.
+DIVA (Costales and Nikolaidis, 2024) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage.
+In sim-to-real settings, GCL (Wang et al., 2024) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
+Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (Narvekar et al., 2020).
+Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (Klink et al., 2020).
+Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (Eimer et al., 2021; Klink et al., 2022).
+WAKER (Rigter et al., 2024) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios.
+Akin to the concept underlying SFL (Rutherford et al., 2024), ProCuRL (Tzannetos et al., 2023) operationalizes the pedagogical notion of the Zone of Proximal Development by using probability of success score-based scoring combined with softmax sampling to prioritize tasks of intermediate difficulty.
 However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
@@
 The agent is trained on these randomly sampled tasks.
-Robust PLR (PLR(REPAIRED) was introduced as an extension of the original PLR (PLR).
+Robust PLR (PLR(Jiang et al., 2021a) was introduced as an extension of the original PLR (Jiang et al., 2021b).
 While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay.
@@
 By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
-Adversarially Compounding Complexity by Editing Levels (ACCEL) (ACCEL) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
+Adversarially Compounding Complexity by Editing Levels (ACCEL) (Parker-Holder et al., 2022) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space.
 This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret.
@@
 Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
-Coverage-based Evaluation of Novelty In Environment (CENIE) (Jayden2024CENIE) introduces environment novelty as a complementary objective to regret in UED.
+Coverage-based Evaluation of Novelty In Environment (CENIE) (Teoh et al., 2024) introduces environment novelty as a complementary objective to regret in UED.
 CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum.
@@
 CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
-Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung2024AdversarialDiffusion) frames environment generation as a diffusion process steered by the agent’s regret.
+Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung et al., 2024) frames environment generation as a diffusion process steered by the agent’s regret.
 Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
-In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (dcd_github), and assessed ADD with the ADD codebase (Chung2024AdversarialDiffusion).
+In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (Jiang et al., 2022), and assessed ADD with the ADD codebase (Chung et al., 2024).
 CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
@@
 While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum.
-Off-policy methods such as SAC (haarnoja2018soft) and TD3 (fujimoto2018addressing) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
-Moreover, Model-based approaches like DreamerV3 (hafner2025) build world models for planning, offering complementary insights into curriculum design.
+Off-policy methods such as SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework.
+Moreover, Model-based approaches like DreamerV3 (Hafner et al., 2025) build world models for planning, offering complementary insights into curriculum design.
 Evaluating these methods under our curriculum framework remains an important direction for future work.
@@
 We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost.
-While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (REPAIRED) and CENIE (Jayden2024CENIE)) and conduct comprehensive experiments to assess its performance in these settings.
+While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (Jiang et al., 2021a) and CENIE (Teoh et al., 2024)) and conduct comprehensive experiments to assess its performance in these settings.
 Evaluation in other domains.
-Our experimental validation has been conducted in MiniGrid (MinigridMiniworld23; PAIRED) and BipedalWalker (Brockman2016Gym; ACCEL).
-Evaluating TRACED in other domains such as Car Racing (Brockman2016Gym; REPAIRED), physical robotics applications (mahmood2018benchmarking), and reasoning task ARC (chollet2019measure) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
+Our experimental validation has been conducted in MiniGrid (Chevalier-Boisvert et al., 2023; Dennis et al., 2020) and BipedalWalker (Brockman et al., 2016; Parker-Holder et al., 2022).
+Evaluating TRACED in other domains such as Car Racing (Brockman et al., 2016; Jiang et al., 2021a), physical robotics applications (Mahmood et al., 2018), and reasoning task ARC (Chollet, 2019) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
 Appendix M Experimental Setup and Reproducibility
@@
 Table 10 summarizes the hyperparameters used.
-Unless otherwise noted, values are adopted directly from the DCD repository (dcd_github).
+Unless otherwise noted, values are adopted directly from the DCD repository (Jiang et al., 2022).
 The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4).
@@
 20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
-Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal2021Reliable.
+Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal et al.
+(2021).
 We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k).
@@
 Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.).
-ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Jayden2024CENIE to our ACCEL implementation.
+ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Teoh et al.
+(2024) to our ACCEL implementation.
 At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.
@@
 Figure 23: TRACED vs. SFL and Robust PLR on multi-agent JaxNav. TRACED consistently outperforms Robust PLR and achieves performance comparable to SFL across all metrics.
-Sampling for Learnability (SFL) (rutherford2024no) is a learnability-driven UED method that prioritizes levels with intermediate difficulty by selecting tasks with high learnability scores, computed as p(1p)p(1-p) where $p$ is the agent’s empirical success rate on a given level obtained from rollouts.
-We evaluate TRACED on the multi-agent JaxNav benchmark following the evaluation protocol introduced in rutherford2024no, including CVaR and mean win rate computed over sampled evaluation levels.
+Sampling for Learnability (SFL) (Rutherford et al., 2024) is a learnability-driven UED method that prioritizes levels with intermediate difficulty by selecting tasks with high learnability scores, computed as p(1p)p(1-p) where $p$ is the agent’s empirical success rate on a given level obtained from rollouts.
+We evaluate TRACED on the multi-agent JaxNav benchmark following the evaluation protocol introduced in Rutherford et al.
+(2024), including CVaR and mean win rate computed over sampled evaluation levels.
 While the original SFL experiments trained baselines for 22,850 PPO updates, we perform our comparison using 4,550 updates for computational efficiency.
-This training regime corresponds to a phase in which SFL is reported to exhibit strong performance relative to other UED baselines (rutherford2024no).
-Aside from the inclusion of TRACED, all training and evaluation details (e.g., level sampling, CVaR estimation, and construction of evaluation sets) match those in rutherford2024no.
+This training regime corresponds to a phase in which SFL is reported to exhibit strong performance relative to other UED baselines (Rutherford et al., 2024).
+Aside from the inclusion of TRACED, all training and evaluation details (e.g., level sampling, CVaR estimation, and construction of evaluation sets) match those in Rutherford et al.
+(2024).
 Results are averaged over three seeds.
 Figure 24 shows CVaR across several risk levels and the mean win rate on the evaluation set.
-In addition to SFL, we include PLR (Robust PLR), the strongest baseline used in rutherford2024no.
+In addition to SFL, we include PLR (Robust PLR), the strongest baseline used in Rutherford et al.
+(2024).
 TRACED consistently outperforms Robust PLR across all CVaR levels and achieves performance comparable to SFL in both CVaR and mean win rate.